### v5

# Holomorphic foliations of degree two and arbitrary dimension

Abstract.
We prove a complete classification for degree-$2$ foliations on $\mathbb{P}^{n}$ of any dimension, provided they are not algebraically integrable. If $\mathscr{F}$ is such a foliation, then either $\mathscr{F}$ is the linear pull-back of a degree 2 foliation by curves on $\mathbb{P}^{n-k+1}$, or a logarithmic foliation of type $(1^{n-k+1},2)$, or a logarithmic foliation of type $(1^{n-k+3})$, or a linear pull-back of a degree-2 foliation of dimension 2 on $\mathbb{P}^{n-k+2}$ tangent to an action of the Lie algebra $\mathfrak{aff}({\mathbb{C}})$. Meanwhile, we prove that any 2-dimensional foliation tangent to a global vector field must satisfy that its tangent sheaf either is not locally free or has a direct summand isomorphic to ${\mathcal{O}}_{\mathbb{P}^{n}}(a)$, $a\in\{0,1\}$. As a byproduct of our classification, we describe the geometry of Poisson structures on $\mathbb{P}^{4}$ with generic rank two.
Key words and phrases:
Holomorphic foliations, Poisson structures, Moduli spaces
2020 Mathematics Subject Classification:
Primary: 37F75, 58A17,14D20,14J60. Secondary:53D17, 14F06
1. Introduction
One of the central problems in the global theory of holomorphic foliations is the description and classification of the irreducible components of the space of foliations on the projective space $\mathbb{P}^{n}$}, $n\geq 3$, at least for low degrees. A complete classification is known for degrees 0 and $1$ in any dimension, and for degree $2$ only in dimension n1n-1. The space $\mathbb{F}{\rm ol}(d,k,n)$ of degree-$d$ dimension-$k$ foliations on $\mathbb{P}^{n}$} is a projective subvariety of ${\mathbb{P}}H^{0}(\Omega_{\mathbb{P}^{n}}^{p}(n-k+d+1))$ whose general element is an integrable (nk)(n-k)-form without zeros of codimension one. A classical result of Darboux implies that every degree-0 foliation is given by the fibers of a linear rational map ${\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}^{n-p}}$, thus $\mathbb{F}{\rm ol}(0,k,n)$ is isomorphic to the Grassmannian $G(k,n+1)$ in its Plücker embedding, see [15, Théorème 3.8]. In degree $d=1$, Loray, Pereira, and Touzet [26], generalizing upon Jouanolou [24], proved that $\mathbb{F}{\rm ol}(1,k,n)$ has two irreducible components: one parameterizing linear pullbacks of degree-one foliations by curves on $\mathbb{P}^{n-k+1}$}; the other parameterizing rational foliations of type $(1^{n-k-1},2)$, given by the fibers of a rational map ${\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}}(1^{n-k-1},2)$. For $d=2$ and $p=1$, Cerveau and Lins Neto [6] proved that $\mathbb{F}{\rm ol}(2,n-1,n)$ has six irreducible components and described their general element. Since then, several works investigated the classification in dimension n1n-1 and higher degrees [7, 17, 28]; most remarkably [12] where the authors present an almost complete classification for degree $3$. In a different direction, the present work focuses on degree-$2$ foliations of dimension $k\leq n-2$. The main result of this work is the following.
Theorem A.
Let $\mathscr{F}$ be a degree-$2$ foliation on $\mathbb{P}^{n}$} of dimension $k\geq 2$. Then one of the following holds:
  1. 1)
    $\mathscr{F}$ is a linear pullback of a rational foliation either of type ${\rm Rat}(1,2,2;n)$ or ${\rm Rat}(1,1,3;n)$;
  2. 2)
    $\mathscr{F}\in{\rm Log}(1^{n-k+1},2;k,n)$ is a logarithmic foliation of type $(1^{n-r+1},2)$;
  3. 3)
    $\mathscr{F}\in{\rm Log}(1^{n+1};k,n)$ is a logarithmic foliation of type $(1^{n+1})$;
  4. 4)
    $\mathscr{F}\in{\rm ALie}(r_{1},\dots,r_{m},k;n)$ is an Affine-Lie foliation; or
  5. 5)
    $\mathscr{F}\in{\rm LPB}(2,k,n)$ is a linear pullback of a foliation by curves in $\mathbb{P}^{n-k+1}$}.
The notation is explained in §2. Theorem A gives a precise description of the irreducible components of $\mathbb{F}{\rm ol}(2,k,n)$ whose general element is not algebraically integrable with either $k>2$ and semistable tangent sheaf; or $k=2$ and stable tangent sheaf. Suppose $\mathscr{F}$ is an algebraically integrable foliation of degree $2$ defined by a rational (nk)(n-k)-form without zeros in codimension one. In that case, it is known that $\mathscr{F}$ belongs to one of the rational components: ${\rm Rat}(1^{n-k-1},2,2;n)$ or ${\rm Rat}(1^{n-k},3;n)$, see Remark 2.1. To decide whether every foliation in item (1) of Theorem A, with $k\leq n-2$, belongs to one of the rational components is a challenging problem whose solution still eludes us. For k=n1k=n-1 it is true, see [26, Proposition 3.5].
The Affine-Lie foliations in item (4) of the theorem come from actions of the Lie algebra of affine transformations of the line $\mathfrak{aff({\mathbb{C}})}$, and these correspond to integer partitions $r_{1}+\dots+r_{m}=n-k+3$. We refer to Pereira-dos Santos [32] for a comprehensive description of such foliations. The number of irreducible components of $\mathbb{F}{\rm ol}(2,k,n)$ grows with the number p(nk+3)p(n-k+3) of partitions of nk+3n-k+3:
$\{\text{irreducible components of }\mathbb{F}{\rm ol}(2,k,n)\}\geq p(n-k+3)+1,$
see [32, Corollary D].
Along our way to proving Theorem A we proved the following result, see Proposition 3.6. We observe that this result was also obtained by da Costa [13] with independent techniques.
Proposition B.
Let $\mathscr{F}$ be a $2$-dimensional foliation of degree $d$ on $\mathbb{P}^{n}$} such that $h^{0}({\rm T}\mathscr{F})>0$. Then, either ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}(a)\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(1+a-d)$, $a\in\{0,1\}$, or $\rm T$$\mathscr{F}$ is not locally free.
Related to holomorphic foliations is the study of Poisson structures, see §5 for a definition. A Poisson structure of generic rank $2$ on $\mathbb{P}^{n}$} corresponds to a foliation of dimension $2$ and degree at most $2$. In this framework, Polishchuk found in [33] a partial classification of foliations of degree $2$ on $\mathbb{P}^{3}$}. Coming from the foliations side, Loray, Pereira, and Touzet [26] classified Poisson structures on Fano $3$-folds of Picard rank one. In [30, section 5.3], Matviichuk, Pym, and Schedler classified Poisson structures on even-dimensional projective spaces which are log symplectic along the divisor $H$ given by the union of the coordinate hyperplanes, i.e., generically symplectic Poisson structures whose the degeneracy locus is the Poisson divisor $H$.
Let $({\mathbb{P}^{n}},\sigma)$ be a holomorphic Poisson structure such that $\sigma\wedge\sigma=0$. Suppose that the associated symplectic foliation $\mathscr{F}_{\sigma}$ has no divisorial zeros. It follows from Theorem A that the tangent sheaf $T\mathscr{F}_{\sigma}$ of the associated symplectic foliation is either split, semi-stable and logarithmic, or stable with algebraic leaves. This means that either the bivector $\sigma$ is globally decomposable or non-decomposable with some Poisson divisor $D$. In general, we will say that a non-decomposable bivector generic rank 2 is of the normal crossing type, if so is $D$.
Due to our study of foliations, we prove the following result about the geometry of Poisson structures with generic rank 2.
Corollary C.
Let $({\mathbb{P}}^{4},\sigma)$ be a general holomorphic Poisson structure such that $\sigma\wedge\sigma=0$. Suppose that the associated symplectic foliation $\mathscr{F}_{\sigma}$ has no divisorial zeros.
If $\sigma$ is decomposable, then either
  • $\sigma=v_{0}\wedge v_{1}\in\wedge^{2}\mathfrak{h}$, where $\mathfrak{h}=H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}(-\log D))$ is the abelian Lie algebra of vector fields tangent to an arrangement of Poisson planes $D=\bigcup_{i=1}^{5}H_{i}$, and the vector fields $v_{i}$ induce on each symplectic leaf of $\sigma$ a transversely Euclidean holomorphic foliation.
  • $\sigma=v_{0}\wedge v$, where $v_{0}$ is a constant vector field, $v$ is a homogeneous polynomial vector field of degree $2$, $v_{0},v]=0$, and $H$ is the Poisson hyperplane determined by $v_{0}\in{\mathbb{P}}H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}(-1))\cong\check{{\mathbb{P}}}^{4}$.
  • $\sigma=v_{0}\wedge v_{1}\in\wedge^{2}\mathfrak{h}$, where $\mathfrak{h}=H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4})$ and $v_{0}$ and $v_{1}$ are linear vector fields such that $v_{0},v_{1}]=v_{0}$, and the vector field $v_{0}$ induces on each symplectic leaf of $\sigma$ a transversely affine holomorphic foliation. Moreover, up to a linear change of homogeneous coordinates $x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]$ on ${\mathbb{P}}^{4}$, the pair $(v_{0},v_{1})$ is of one of the following two types:
    • (I)
      Type $(4,1)$.
      $v_{0}=x_{0}\,\frac{\partial}{\partial x_{1}}+x_{1}\,\frac{\partial}{\partial x_{2}}+x_{2}\,\frac{\partial}{\partial x_{3}},$
      $v_{1}=x_{1}\,\frac{\partial}{\partial x_{1}}+2x_{2}\,\frac{\partial}{\partial x_{2}}+3x_{3}\,\frac{\partial}{\partial x_{3}},$
      so that $v_{0},v_{1}]=v_{0}$ and the Jordan type of $v_{0}$ is $(4,1)$.
    • (II)
      Type $(3,2)$.
      $v_{0}=x_{0}\,\frac{\partial}{\partial x_{1}}+x_{1}\,\frac{\partial}{\partial x_{2}}+x_{3}\,\frac{\partial}{\partial x_{4}},$
      $v_{1}=x_{1}\,\frac{\partial}{\partial x_{1}}+2x_{2}\,\frac{\partial}{\partial x_{2}}+x_{4}\,\frac{\partial}{\partial x_{4}},$
      so that $v_{0},v_{1}]=v_{0}$ and the Jordan type of $v_{0}$ is $(3,2)$.
If $\sigma$ is not decomposable and of normal crossing type, then either
  • $\sigma\in H^{0}\bigl({\mathbb{P}}^{4},\wedge^{2}T_{{\mathbb{P}}^{4}/{\mathbb{P}}^{2}}(-\log D)\bigr)$, where $D=H_{1}\cup H_{2}\cup\mathcal{V}$ is a Poisson divisor such that $H_{1}$ and $H_{2}$ are hyperplanes and $\mathcal{V}$ is a cubic, or
  • $\sigma\in H^{0}\bigl({\mathbb{P}}^{4},\wedge^{2}T_{{\mathbb{P}}^{4}/{\mathbb{P}}^{2}}(-\log D)\bigr)$, where $D=H_{1}\cup\mathcal{Q}_{1}\cup\mathcal{Q}_{2}$ is a Poisson divisor such that $H_{1}$ is a hyperplane and $\mathcal{Q}_{1},\mathcal{Q}_{2}$ are quadrics, or
  • $\sigma\in H^{0}\bigl({\mathbb{P}}^{4},\wedge^{2}T{\mathbb{P}}^{4}(-\log D)\bigr)$, where $D=H_{1}\cup H_{2}\cup H_{3}\cup\mathcal{Q}$ is a Poisson divisor such that $H_{i}$ are hyperplanes and $\mathcal{Q}$ is a quadric, and there exists $v\in H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}(-\log D))$ which is Hamiltonian with respect to $\sigma$.
Consider the space of Poisson structures of generic rank 2
$\mathrm{Poisson}(X)_{2}:=\left\{\sigma\in\mathbb{P}H^{0}(X,\wedge^{2}TX)\,\Big|\,[\sigma,\sigma]=0,\sigma\wedge\sigma=0\,\right\}\subset\mathrm{Poisson}(X).$
As a consequence of the study of foliations of degree 2 and dimension 2 we obtain.
Corollary D.
The space $\mathrm{Poisson}(\mathbb{P}^{4})_{2}$ has at least seven irreducible components.
This work is organized as follows. In Section 2 we establish notation and recall basic results on holomorphic foliations. We describe in §§2.2 the $2$-dimensional foliations associated with an action of $\mathfrak{aff({\mathbb{C}})}$, see Theorem 2.3. We call them Affine-Lie foliations. We also describe a rational first integral for a general foliation associated with the partition $(n+1)$, see Proposition 2.4. In Section 3, we give a detailed study of subfoliations, especially those arising by strict semistability. In particular, we prove Proposition B (see Proposition 3.6) and prepare for the proof of Theorem A. In Section 4, we prove Theorem A by proving two results: Theorem 4.1 for 2-dimensional foliations, and Theorem 4.2 for higher dimensions. Finally, in Section 5, we discuss Poisson structures and we prove Corollary C.
Acknowledgments
We thank Jorge Vitório Pereira for his interest in this work and for numerous insightful conversations. MC is partially supported by the Università degli Studi di Bari and by the PRIN 2022MWPMAB- "Interactions between Geometric Structures and Function Theories" and is a member of INdAM-GNSAGA. AM was partially supported by INCTmat/MCT/Brazil, CNPq grant number 160934/2022-2.
2. Holomorphic foliations
In this section, we review the basic concepts needed throughout the work. Let $X$ be a smooth complex projective variety; an experienced reader may find no difficulty relaxing this requirement. A singular holomorphic distribution $\mathscr{F}$ on $X$ is the data of a saturated subsheaf $\rm T$$\mathscr{F}$ of the tangent sheaf ${\rm T}X$ of $X$. These sheaves give a short exact sequence
(2.1) $0\longrightarrow{\rm T}\mathscr{F}\longrightarrow{\rm T}X\longrightarrow{\rm N}\mathscr{F}\longrightarrow 0$
where $\rm N$$\mathscr{F}$ is a torsion-free sheaf, called the normal sheaf of $\mathscr{F}$, and $\rm T$$\mathscr{F}$ is a reflexive sheaf, called the tangent sheaf of $\mathscr{F}$. The dimension of $\mathscr{F}$ is the generic rank of $\rm T$$\mathscr{F}$. If $\rm T$$\mathscr{F}$ is involutive, i.e. ${\rm T}\mathscr{F},{\rm T}\mathscr{F}]\subset{\rm T}\mathscr{F}$, then $\mathscr{F}$ is a foliation; this means that through a general point in $X$ passes a unique connected immersed (analytic) subvariety $L\subset X$ such that ${\rm T}L={\rm T}\mathscr{F}|_{L}$, called a leaf. A foliation $\mathscr{F}$ is algebraically integrable if its leaves are embedded algebraic varieties.
Taking exterior power on (2.1) we get a morphism $\omega_{\mathscr{F}}\colon{\textstyle\bigwedge}^{p}{\rm T}X\to\det({\rm N}\mathscr{F})$, where $p=\operatorname{rk}{\rm N}\mathscr{F}$, whose image is $\det({\rm N}\mathscr{F})$ tensored by the ideal sheaf $\mathscr{I}_{Z}$ of the vanishing locus of $\omega_{\mathscr{F}}$ seen as a global section of $\Omega_{X}^{p}\otimes\det({\rm N}\mathscr{F})$. We call $Z$ the singular scheme of $\mathscr{F}$, and denote it by $\operatorname{Sing}(\mathscr{F})$. Since $\rm T$$\mathscr{F}$ is saturated, $\operatorname{codim}(\operatorname{Sing}(\mathscr{F}))\geq 2$.
Given a distribution $\mathscr{F}$, the twisted $p$-form $\omega_{\mathscr{F}}\in H^{0}(\Omega_{X}^{p}\otimes\det({\rm N}\mathscr{F}))$ is rather special, for any $p\in X\setminus\operatorname{Sing}(\mathscr{F})$ there exists an open neighborhood $U\ni p$ and $1$-forms $\eta_{1},\dots,\eta_{p}\in\Omega_{X}^{1}(U)$ such that
$\omega_{\mathscr{F}}|_{U}=\eta_{1}\wedge\dots\wedge\eta_{p}.$
Moreover, $\mathscr{F}$ is a foliation if and only if
$d\eta_{j}\wedge\omega_{\mathscr{F}}|_{U}=0,\,\text{for}\,j=1,\dots,p.$
2.1. Foliations on projective spaces
A $k$-dimensional distribution $\mathscr{F}$ on $\mathbb{P}^{n}$} corresponds to a twisted (nk)(n-k)-form $\omega\in H^{0}(\Omega_{{\mathbb{P}^{n}}}^{n-k}(d+n-k+1)$, where $d\geq 0$ is called the degree of $\mathscr{F}$. The Euler sequence allows us to interpret $\omega$ as a polynomial differential form
$\omega=\sum_{i=0}^{n}A_{i}dx_{i}$
where $A_{i}$ are homogeneous of degree $d+1$, and the contraction with the radial vector files $\operatorname{rad}=x_{0}\frac{\partial}{\partial x_{0}}+\cdots+x_{n}\frac{\partial}{\partial x_{n}}$ vanishes: $\iota_{\operatorname{rad}}\omega=0$. The local decomposability translates to the Plücker conditions
(2.2) $({i}_{v}\omega)\wedge\omega=0\quad\text{ for every }v\in\wedge^{n-k-1}(\mathbb{C}^{n+1}),$
and the integrability condition becomes
(2.3) $({i}_{v}\omega)\wedge d\omega=0\quad\text{ for every }v\in\wedge^{n-k-1}(\mathbb{C}^{n+1}).$
We refer to [24, 10] for details.
Given integers $d$, $k$, and $n$, the conditions (2.2) and (2.3) define the following subvariety of ${\mathbb{P}}H^{0}(\Omega_{{\mathbb{P}^{n}}}^{n-k}(n-k+d+1))$:
(2.4) $\mathbb{F}{\rm ol}(d,k,n)=\Big\{[\omega]\mid\omega\text{ is integrable, and }\operatorname{codim}(\operatorname{Sing}\omega)\geq 2\Big\},$
the space of $k$-dimensional foliations of degree $d$ on $\mathbb{P}^{n}$}. For $d\leq 1$ the irreducible components of $\mathbb{F}{\rm ol}(d,k,n)$ are well-known. A degree-0 foliation is defined by a linear rational map ${\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}^{n-k}}$, see [15, Théorème 3.8]. In particular,
$\mathbb{F}{\rm ol}(0,n-k,n)\cong G(n-k+1,n+1)$
the Grassmannian of (nk+1)(n-k+1)-planes in ${\mathbb{C}}^{n+1}$. For degree one, there is the following classification.
Theorem 2.1.
[26, Theorem 6.2] If $\mathscr{F}$ is a foliation of degree $1$ and codimension $q$ on $\mathbb{P}^{n}$}, then we are in one of the following cases:
  1. 1)
    $\mathscr{F}$ is defined by a dominant rational map $\rho\colon{\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}}(1^{(q+1)},2)$, with irreducible general fiber determined by $q$ linear forms and one quadratic form; or
  2. 2)
    $\mathscr{F}$ is the linear pullback of a foliation induced by a global holomorphic vector field on $\mathbb{P}^{n-k+1}$.
Hence, $\mathbb{F}{\rm ol}(1,n-k,n)$ has two irreducible components. In degree $2$, only the case of codimension-one foliations is known.
Theorem 2.2.
[6, Theorem A] The space of codimension-one foliations of degree $2$ in $\mathbb{P}^{n}$}, $n\geq 3$, has $6$ irreducible components:
  • S(2,n){S(2,n)} composed of linear pullbacks of degree-$2$ foliations on $\mathbb{P}^{2}$};
  • R(2,2){R(2,2)} given by pencils of quadrics;
  • R(1,3){R(1,3)} given by pencils of cubics with a triple hyperplane;
  • L(1,1,1,1){L(1,1,1,1)} whose general member is defined by a logarithmic $1$-form with poles in $4$ hyperplanes;
  • L(1,1,2){L(1,1,2)} whose general member is defined by a logarithmic $1$-form with poles in two hyperplanes and a quadric;
  • E(n){E(n)} the exceptional component whose general member is defined by an algebraic action of the Lie algebra of affine transformations $\mathfrak{aff}({\mathbb{C}})$.
Below, we detail each one of these components. We describe their generalizations to any codimension.
Example 2.1 (Linear pullback foliations).
Given a linear projection $\rho\colon{\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}^{n-k+1}}$ and a $\mathscr{G}$ degree-$d$ foliation by curves on $\mathbb{P}^{n-k+1}$}, then we define $\mathscr{F}=\rho^{*}\mathscr{G}$ a degree-$d$ foliation of dimension $k$. Varying $\rho$ and $\mathscr{G}$, we get an irreducible component ${\rm LPB}(d,k,n)\subset\mathbb{F}{\rm ol}(d,k,n)$. For a proof, see [11, Corollary 5.1].
Example 2.2 (Rational foliations).
Given homogeneous polynomials $F_{0},\dots,F_{n-k}$, we have a rational map $\rho\colon{\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}}(d_{0},\dots,d_{n-k})$, where $d_{j}=\deg(F_{j})$. Its fibers define an algebraically integrable foliation that we call rational if $\rho$ does not ramify in codimension one, i.e., the maximal minors of the Jacobian matrix of $\rho$ do not share a common factor. These foliations fill an irreducible component
${\rm Rat}(d_{0},\dots,d_{n-k};n)\subset\mathbb{F}{\rm ol}(d,k,n),$
where $d=-n+k-1+\sum_{j=0}^{n-k}d_{j}$. For a proof, see [10].
Remark 2.1.
In degree $2$, there are two components of rational foliations. Indeed, the equation $2=-n+k-1+\sum_{j=0}^{n-k}d_{j}$ admits only two solutions:
$(d_{0},\dots,d_{n-k})=(1^{n-k-2},2,2)\text{ or }(1^{n-k-1},3).$
If a foliation $\mathscr{F}$ of dimension $k$ and degree 2 is tangent to a rational map $\rho\colon{\mathbb{P}^{n}}\dashrightarrow\mathbb{P}(d_{0},\dots,d_{n-k})$, such that the $\mathscr{F}$-invariant divisor $\{F_{0}=\dots=F_{n-k}=0\}$ is normal crossing, then by [4] we get $d_{0}+\cdots+d_{n-k}\leq n-k+3$. This implies that $\mathscr{F}$ is rational.
Example 2.3 (Logarithmic foliations).
A logarithmic foliation on $\mathbb{P}^{n}$} of dimension $k$ is defined by a closed rational (nk)(n-k)-form
$\omega=\sum_{0\leq i_{1}<\cdots<i_{n-k}\leq r}\lambda_{i_{1},\dots,i_{n-k}}\frac{dF_{i_{1}}}{F_{i_{1}}}\wedge\cdots\wedge\frac{dF_{i_{n-k}}}{F_{i_{n-k}}}$
where $F_{0},\dots,F_{r}$ are homogeneous polynomials of degrees $d_{0},\dots,d_{r}$, with $r\geq n-k$, and $\lambda_{i_{1},\dots,i_{n-k}}\in{\mathbb{C}}$ satisfy Plücker conditions. We also assume that $\omega$ does not vanish in codimension one. Owing to [8, Theorem 1.17], if $r\geq n-k+1$ these foliations form an irreducible component
${\rm Log}(d_{0},\dots,d_{r};k,n)\subset\mathbb{F}{\rm ol}(d,k,n),$
where $d=-n+k-1+\sum_{j=0}^{r}d_{j}$. We remark that for r=nkr=n-k these are rational foliations.
Remark 2.2.
In degree $2$, the equation $2=-n+k-1+\sum_{j=0}^{r}d_{j}$ has only two solutions with $r\geq n-k+1$:
$(d_{0},\dots,d_{r})=(1^{n-k+1},2)\text{ or }(1^{n-k+3}).$
Thus $\mathbb{F}{\rm ol}(2,k,n)$ has two components of logarithmic type. We observe that a foliation of dimension $k$ is a general logarithmic foliation of type $(1^{n-k+1},2)$ if and only if it is the pullback, via a rational map $\rho\colon{\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}}(1^{n-k+1},2)$, of a foliation defined by a global vector field. The proof is a straightforward computation that we leave to the reader.
A logarithmic foliation of type $(1^{n-k+3})$ and dimension $k>2$ is a linear pullback of a $2$-dimensional logarithmic foliation of the same type. Indeed, for $k>2$, nk+3<n+1n-k+3<n+1, and, up to a linear change of coordinates, the foliation is defined by fewer than $n+1$ variables. For dimension $k=2$, a general logarithmic foliation of type $(1^{n+1})$ is defined by a pair commuting global vector fields on $\mathbb{P}^{n}$}, i.e., it is associated with an action of the commutative Lie algebra ${\mathbb{C}}^{2}\hookrightarrow H^{0}(T{\mathbb{P}^{n}})$.
2.2. Foliations associated with a Lie algebra action
Given a $k$-dimensional Lie subalgebra $\mathfrak{g}\subset\mathfrak{sl}_{n+1}=H^{0}(T{\mathbb{P}^{n}})$, it defines an element $\omega(\mathfrak{g})\in H^{0}(\Omega_{{\mathbb{P}^{n}}}^{n-k}(n+1))$. If $\omega(\mathfrak{g})\neq 0$ does not vanish in codimension one, it defines a foliation of degree equal to its dimension, $d=k$. We refer to [15, 11, 32] for generalities and details about these foliations. Here we will focus on $2$-dimensional Lie algebras. In [32], Pereira and dos Santos describe the scheme of Lie subalgebras and the irreducible components of the space of foliations they define.
Consider the algebraic variety of $2$-dimensional Lie subalgebras of $\mathfrak{sl}_{n+1}$:
${\rm SLie}(2,\mathfrak{sl}_{n+1})=\{\,V\subset\mathfrak{sl}_{n+1}\mid[V,V]\subset V,\,\dim V=2\,\}.$
It is a closed subvariety of the Grassmannian $G(2,\mathfrak{sl}_{n+1})$ defined by the equations imposed by the Lie bracket. ${\rm SLie}(2,\mathfrak{sl}_{n+1})$ has several irreducible components, one $\mathcal{C}$ corresponding to commutative Lie algebras and another for each integer partition of $n+1$:
${\rm SLie}(2,\mathfrak{sl}_{n+1})=\mathcal{C}\cup\bigcup_{r_{1}+\cdots+r_{m}=n+1}\mathcal{S}_{(r_{1},\dots,r_{m})}$
Consider the map $\mathcal{F}\colon{\rm SLie}(2,\mathfrak{sl}_{n+1})\longrightarrow\mathbb{F}{\rm ol}(2,n-2,n)$ associating with a Lie subalgebra $\mathfrak{g}$ the foliation $\mathcal{F}(\mathfrak{g})$ it defines, whenever it makes sense. We saw in Remark 2.2 that $\mathcal{F}(\mathcal{C})={\rm Log}(1^{n+1};2,n)$ the logarithmic component. For the other components, there is the following result, which is a restatement of [32, Theorem 3.6]. They only consider partitions with one part at least $5$; however, the argument works for other partitions with minor changes.
Theorem 2.3.
[32, Theorem 3.6] We have that ${\rm ALie}(r_{1},\dots,r_{m},2;n):=\mathcal{F}(\mathcal{S}_{(r_{1},\dots,r_{m})})$ is an irreducible component of $\mathbb{F}{\rm ol}(2,n-2,n)$ unless $(r_{1},\dots,r_{m})$ is one of the following partitions:
$\{(3,1^{n-2}),(2,2,1^{n-3}),(2,1^{n-1}),(1^{n+1})\}.$
Moreover, $\mathcal{F}$ is not defined on $\mathcal{S}_{(1^{n+1})}$ and $\mathcal{S}_{(2,1^{n-1})}$.
Remark 2.3.
The partition $(4)$ gives the classical exceptional component on $\mathbb{P}^{3}$} [6]. The partition $(n+1)$ for $n\geq 3$ figures in [11, Example 6.8]. These are all rigid, i.e., the component ${\rm ALie}(n+1,2;n)$ is the orbit closure of a general foliation in it. We also remark that $\mathcal{F}(\mathcal{S}_{(3,1^{n-2})})$ and $\mathcal{F}(\mathcal{S}_{(2,2,1^{n-3})})$ do not fill irreducible components of $\mathbb{F}{\rm ol}(2,n-2,n)$, both fall into the logarithmic component ${\rm Log}(1^{n-1},2;2,n)$, see Proposition 3.9 below.
Remark 2.4.
Due to [11, Corollary 5.1], if $k\geq 3$ and $r_{1}+\dots+r_{m}=n-k+3$, there are irreducible components ${\rm ALie}(r_{1},\dots,r_{m},k;n)\subset\mathbb{F}{\rm ol}(2,k,n)$ whose general element is a linear pullback of a foliation in ${\rm ALie}(r_{1},\dots,r_{m},2;n-k+2)$. We call these the Affine-Lie foliations.
2.2.1. First integrals for Affine-Lie foliations
Fix a partition $r_{1}\geq r_{2}\geq\dots\geq r_{m}$ and fix adapted homogeneous coordinates:
$(x_{0}^{1}:\dots:x_{r_{1}-1}^{1}:\dots:x_{0}^{k}:\dots:x_{r_{m}-1}^{k}).$
Then define the vector fields $v_{(r_{1},\dots,r_{m})}=\delta_{r_{1}}+\cdots+\delta_{r_{m}}$ and $w_{(r_{1},\dots,r_{m})}=\rho_{r_{1}}+\cdots+\rho_{r_{j}}$, where
(2.5) $\delta_{r_{j}}=\sum_{i=0}^{r_{j}-2}x_{i}^{j}\frac{\partial}{\partial x_{i+1}^{j}},\quad\text{and}\quad\rho_{r_{j}}=\sum_{i=0}^{r_{j}-1}(r_{j}-1-2i+\epsilon_{j})x_{i}^{j}\frac{\partial}{\partial x_{i}^{j}},\quad\epsilon_{1}+\dots+\epsilon_{k}=0\in{\mathbb{C}}.$
It follows that $w_{(r_{1},\dots,r_{m})},v_{(r_{1},\dots,r_{m})}]=v_{(r_{1},\dots,r_{m})}$. Then $v_{(r_{1},\dots,r_{m})}$ and $w_{(r_{1},\dots,r_{m})}$ define a foliation in ${\rm ALie}(r_{1},\dots,r_{m},2;n)$. We will focus on the partition $(n+1)$.
The vector field $\delta_{m}$, as defined in (2.5), is sometimes called the basic Weitzenböck derivation. In general, a Weitzenböck derivation is a linear nilpotent derivation. They play an important role in Invariant Theory, and the name comes from Weitzenböck’s Theorem (see [35]) which asserts that any linear action of the additive group $\mathbb{G}_{a}$ on a polynomial ring has a finitely generated ring of invariants. That is, the ring
$\ker\delta=\{\,f\in k[y_{1},\dots,y_{n}]\mid\delta(f)=0\,\}$
is finitely generated (over $k$) for $\delta$ a Weitzenböck derivation. This fails for polynomial locally nilpotent derivations of higher degree, see [31, Example 6.3.1]. In [14], the authors give explicit generators of the field of fractions of the constant ring of the basic Weitzenböck derivation. Using this description, we can show that a foliation in ${\rm ALie}(n+1,2;n)$ is algebraically integrable.
Proposition 2.4.
A general Affine-Lie foliation in ${\rm ALie}(n+1,2;n)$ is algebraically integrable.
Proof.
It is enough to prove that the foliation $\mathscr{F}$ defined by $\delta_{n+1}$ and $\rho_{n+1}$ in (2.5) is algebraically integrable. Note that we only need to prove that there exist n2n-2 algebraically independent functions $f_{i}/g_{i}\in{\mathbb{C}}(x_{0},\dots,x_{n})$ such that $f_{i}$ and $g_{i}$ are homogeneous of the same degree, and $v(f_{i}/g_{i})=w(f_{i}/g_{i})=0$.
According to [14, §5], the field of fractions of $\ker v$ is generated by algebraically independent elements $z_{1},\dots,z_{n}$ given by:
$z_{1}$ $=x_{0},$
$z_{2i}$ $=\sum_{j=0}^{2i}(-1)^{j}x_{j}x_{2i-j},\,\text{and}$
$z_{2i+1}$ $=\frac{x_{0}}{2n-4i}\left(\sum_{j=0}^{2i}(-1)^{j}(j+1)(n-j)x_{j+1}x_{2i-j}\right$.
$+\sum_{j=0}^{2i}(-1)^{j}(2i+1-j)(n-2i+j)x_{j}x_{2i+1-j}\Bigg)-x_{1}\sum_{j=0}^{2i}(-1)^{j}x_{j}x_{2i-j},\,\text{for}\,j\geq 1.$
One may verify that $w(z_{1})=\frac{n}{2}z_{1}$, $w(z_{2i})=n-2i$, and $w(z_{2i+1})=\frac{3n}{2}-2i-1$ for $j\geq 1$. A first integral for $\mathscr{F}$ is then a rational function $f=z_{1}^{a_{1}}\cdots z_{n}^{a_{n}}$ such that $w(f)=0$, hence, the exponents $a_{j}$ satisfy
$\begin{cases}\displaystyle a_{1}+\sum_{i=1}^{\lfloor n/2\rfloor}(2a_{2i}+3a_{2i+1})=0\\ \displaystyle\frac{n}{2}a_{1}+\sum_{i=1}^{\lfloor n/2\rfloor}\left((n-2i)a_{2j}+\left(\frac{3n}{2}-2i-1\right)a_{2j+1}\right)=0\end{cases}.$
The $2\times n$ matrix $A$ associated with this system is of rank $2$. Then we can find n2n-2 integer-valued solutions, linearly independent over $\mathbb{Q}$, which provide n2n-2 algebraically independent rational first integrals for $\mathscr{F}$. Explicitly, an independent set of first integrals is:
$\left\{\frac{z_{2+i}^{2}z_{1}^{4\lfloor i/2\rfloor}}{z_{2}^{i+2}},\,i=1,\dots,n-2\right\}.$
for $i$ even one can take a square root. ∎
From the above proof, we also conclude that $\mathscr{F}$ is given by the (n2)(n-2)-form
$\omega=\frac{1}{z_{1}^{\frac{n^{2}-n-4}{2}}}\iota_{\operatorname{rad}}dz_{2}\wedge dz_{3}\wedge d(z_{1}^{2}z_{4})\wedge d(z_{1}^{2}z_{5})\wedge d(z_{1}^{4}z_{6})\wedge\dots\wedge d(z_{1}^{2\lfloor\frac{n-2}{2}\rfloor}z_{n}).$
3. Study of subfoliations
In this section, we study subfoliations $\mathscr{G}\subset\mathscr{F}$ of a fixed foliation $\mathscr{F}$, with special attention to subfoliations defined by actions of Lie algebras. A Lie algebra naturally attached to a foliation $\mathscr{F}$ on $\mathbb{P}^{n}$} is given by its infinitesimal automorphisms
$\mathfrak{aut}(\mathscr{F})=\{\,v\in H^{0}(T{\mathbb{P}^{n}})\mid\exists c\in{\mathbb{C}}\text{ s.t. }\mathcal{L}_{v}\omega=c\,\omega\,\},$
where $\mathcal{L}_{v}$ stands for the Lie derivative in the direction of $v$ and $\omega$ is a twisted differential form defining $\mathscr{F}$. $\mathfrak{aut}(\mathscr{F})$ is the Lie algebra of the closed algebraic subgroup $\operatorname{Aut}(\mathscr{F})\subset\operatorname{PGL}(n+1,{\mathbb{C}})$ that preserves $\mathscr{F}$. The global vector fields tangent to $\mathscr{F}$ belong to $H^{0}({\rm T}\mathscr{F})$ and define a Lie subalgebra of $\mathfrak{aut}(\mathscr{F})$, usually denoted by $\mathfrak{fix}(\mathscr{F})$. The exponential of $\mathfrak{fix}(\mathscr{F})$ is not necessarily a closed subgroup. Let $\mathfrak{afix}(\mathscr{F})$ denote the Lie algebra of the smallest algebraic subgroup whose Lie algebra contains $\mathfrak{fix}(\mathscr{F})$. Note that
$\mathfrak{fix}(\mathscr{F})\subset\mathfrak{afix}(\mathscr{F})\subset\mathfrak{aut}(\mathscr{F}).$
The computation of $\mathfrak{afix}(\mathscr{F})$ may be carried out as in [3, Chapter II, §7].
Another way to obtain special subdistributions/subfoliations is slope-stability. Given a coherent sheaf $F$ on $\mathbb{P}^{n}$} of positive rank $\operatorname{rk}F$, the slope is defined by $\mu(F):=c_{1}(F)/\operatorname{rk}F$. Then $F$ is called semistable if $\mu(F)\geq\mu(E)$ for every $E\subset F$ such that $\operatorname{rk}E<\operatorname{rk}F$. $F$ is called stable if the inequality is always strict. One also defines
$\mu_{\rm max}(F)$ $=\max\{\,\mu(E)\mid E\subset F\,\};$
$\mu_{\rm min}(F)$ $=\min\{\,\mu(Q)\mid F\twoheadrightarrow Q\,\text{and }Q\text{ is torsion free}\,\}.$
We start with the following general observation that we will apply to $\mathfrak{fix}(\mathscr{F})$. Note that the condition $\mu_{\rm max}(F)\leq 0$ holds for $F$ semistable with $c_{1}(F)=0$.
Lemma 3.1.
Let $F$ be a rank-$r$ torsion-free sheaf on $\mathbb{P}^{n}$} such that $\mu_{\rm max}(F)\leq 0$. Then the evaluation morphism is injective
$H^{0}(F)\otimes{\mathcal{O}}_{\mathbb{P}^{n}}\hookrightarrow F$
In particular, $h^{0}(F)\leq r$. If equality holds, then $F={\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus r}$.
Proof.
Let $h=h^{0}(F)$ and consider the evaluation map $\phi\colon{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus h}\to F$ generated by linear independent global sections. We get a short exact sequence
$0\longrightarrow\ker\phi\longrightarrow{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus h}\longrightarrow\operatorname{im}\phi\longrightarrow 0.$
Note that $\operatorname{im}\phi$ must be torsion free. Hence, $\ker\phi\subset{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus h}$ is saturated. Since ${\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus h}$ is semistable, $\mu(\ker\phi)\leq 0$ and $\mu(\operatorname{im}\phi)\geq 0$. On the other hand, $\mu(\operatorname{im}\phi)\leq\mu_{\rm max}(F)\leq 0$. Then $\mu(\operatorname{im}\phi)=\mu(\ker\phi)=0$. By [23, Corollary 1.6.11], $\ker\phi$ must be a direct summand of ${\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus h}$. One can interpret the map $\ker\phi={\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus a}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus h}$ as $\mathbb{C}$}-linear relations among the global sections defining $\phi$. Therefore, $\ker\phi=0$, i.e., $\phi$ is injective, and $h\leq\operatorname{rk}F$.
Suppose $h=\operatorname{rk}F$ and consider the exact sequence
$0\longrightarrow{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus\operatorname{rk}F}\longrightarrow F\longrightarrow\operatorname{coker}\phi\longrightarrow 0.$
Then $\operatorname{coker}\phi$ is a torsion sheaf whose first Chern class is a positive multiple of the divisorial part of its support. On the other hand, $\mu(F)\leq\mu_{\rm max}(F)\leq 0$. Then $c_{1}(Q)=c_{1}(F)\leq 0$, which implies that $Q$ is supported in codimension at least 2. Since $F$ is reflexive, $\phi$ must be an isomorphism. ∎
Proposition 3.2.
Let $\mathscr{F}$ be a foliation on $\mathbb{P}^{n}$} such that $\mu_{\rm max}({\rm T}\mathscr{F})\leq 0$, then it admits a subfoliation of dimension $h^{0}({\rm T}\mathscr{F})$ generated by global vector fields.
Proof.
By Lemma 3.1, the evaluation map $\phi\colon H^{0}({\rm T}\mathscr{F})\otimes{\mathcal{O}}_{\mathbb{P}^{n}}\to{\rm T}\mathscr{F}$ is injective. Then $H^{0}({\rm T}\mathscr{F})\otimes{\mathcal{O}}_{\mathbb{P}^{n}}\hookrightarrow{\rm T}\mathscr{F}\hookrightarrow T{\mathbb{P}^{n}}$ defines a subdistribution, which is a foliation since $H^{0}({\rm T}\mathscr{F})=\mathfrak{fix}(\mathscr{F})$ is closed under the Lie bracket. ∎
Next, we provide a simple technical observation that will be useful to constrain the singular locus of a subfoliation $\mathscr{G}$ whose tangent sheaf is locally free. Exact sequences, such as the one below, are known in Commutative Algebra as Bourbaki sequences; see, for instance, [22] and the references therein.
Lemma 3.3.
Consider an exact sequence of sheaves on $\mathbb{P}^{n}$} (or any smooth variety)
$0\longrightarrow E\longrightarrow F\longrightarrow\mathscr{I}_{W}(a)\longrightarrow 0$
where $E$ is locally free, $F$ is reflexive, and $\operatorname{codim}W\geq 2$. Then either
  • $W=\emptyset$ (hence $F$ is locally free), or
  • $W$ is pure of codimension two and Cohen-Macaulay off $\operatorname{Sing}(F)$.
Proof.
Since $F$ is reflexive, it is locally free off $Z=\operatorname{Sing}(F)$ whose codimension is at least three. Note that $Z\subset W$ (set-theoretically) since $F$ must be locally free at any point off $W$. At any point $p\in W\setminus Z$, the Hilbert-Burch theorem says that $W$ is Cohen-Macaulay of codimension two.
We only need to prove that $W$ is pure or, equivalently, that $\mathcal{O}_{W}$ satisfies the Serre condition $S_{1,2}$, see [23, Proposition 1.1.10]. This condition means that $\operatorname{depth}(\mathcal{O}_{W,p})\geq 1$ for any point $p\in W$ of codimension at least 33. Fix such a point pp. Since $F$ is reflexive and $E$ is locally free, we have $\operatorname{depth}(F_{p})\geq 2$ and $\operatorname{depth}(E_{p})=\operatorname{codim}p\geq 3$. Then $\operatorname{depth}(\mathscr{I}_{W,p})\geq 2$ and $\operatorname{depth}(\mathcal{O}_{W,p})\geq 1$, see [5, Proposition 1.2.9]. ∎
Proposition 3.4.
Let $\mathscr{F}$ be a distribution on $\mathbb{P}^{n}$} such that there exists a subdistribution $\mathscr{G}\subset\mathscr{F}$ satisfying: $\dim\mathscr{G}=\dim\mathscr{F}-1$, ${\rm T}\mathscr{G}$ is locally free and $H^{1}({\rm T}\mathscr{G}(c_{1}({\rm T}\mathscr{G})-c_{1}({\rm T}\mathscr{F})))=0$. Then either
  • ${\rm T}\mathscr{F}\cong{\rm T}\mathscr{G}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(c_{1}({\rm T}\mathscr{F})-c_{1}({\rm T}\mathscr{G}))$, or
  • ${\rm T}\mathscr{F}/{\rm T}\mathscr{G}=\mathscr{I}_{W}(c_{1}({\rm T}\mathscr{F})-c_{1}({\rm T}\mathscr{G}))$ with $W\subset\operatorname{Sing}(\mathscr{G})$ a subscheme of pure codimension two.
Proof.
By hypothesis, $\rm T$$\mathscr{F}$ and ${\rm T}\mathscr{G}$ fit in the following commutative diagram with exact rows and columns, where $\delta=c_{1}({\rm T}\mathscr{F})-c_{1}({\rm T}\mathscr{G})$.
(3.1) ${\rm T}\mathscr{G}$${\rm T}\mathscr{G}$$\rm T$$\mathscr{F}$$T{\mathbb{P}^{n}}$$\rm N$$\mathscr{F}$$\mathscr{I}_{W}(\delta)$${\rm N}\mathscr{G}$$\rm N$$\mathscr{F}$$\scriptstyle{\rho}$$\scriptstyle{\phi}$$\scriptstyle{\psi}$
By Lemma 3.3, either $W=\emptyset$ or $W$ is pure of codimension two. If $W=\emptyset$, then $\rm T$$\mathscr{F}$ corresponds to an extension class in $\operatorname{Ext}^{1}({\mathcal{O}}_{\mathbb{P}^{n}}(\delta),{\rm T}\mathscr{G})=H^{1}({\rm T}\mathscr{G}(-\delta))$ which is zero by hypothesis. Thus ${\rm T}\mathscr{F}={\rm T}\mathscr{G}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(\delta)$ in this case.
Assume $W\neq\emptyset$. By Lemma 3.3, $W$ is pure of codimension two. Let $p$ be an associated point of $\mathcal{O}_{W}$. We have that $p$ has codimension two and, on the other hand, $\rm T$$\mathscr{F}$ is reflexive. Thus ${\rm T}\mathscr{F}_{p}$ is a free $\mathcal{O}_{{\mathbb{P}^{n}},p}$-module. Of course, the same holds for ${\rm T}\mathscr{G}$ and $T{\mathbb{P}^{n}}$. Then, we may regard $\phi_{p}=\psi_{p}\rho_{p}$ as a product of matrices.
The ideal sheaf $\mathscr{I}_{Z}$ of $Z=\operatorname{Sing}(\mathscr{G})$ is defined by the $s\times s$ minors of $\phi$, where $s=\dim\mathscr{G}$. Also note that $\mathscr{I}_{W,p}$ is defined by the $s\times s$ minors of $\rho_{p}$. Then, by Binet-Cauchy Theorem, [19, p.11], the $s\times s$ minors of a product of matrices is a combination of $s\times s$ minors of both matrices. Hence, $\mathscr{I}_{Z,p}\subset\mathscr{I}_{W,p}$. Since $p$ is arbitrary, we get $\mathscr{I}_{Z}\subset\mathscr{I}_{W}$, see [29, §6]. Therefore, $W\subset\operatorname{Sing}\mathscr{G}$. ∎
Remark 3.1.
1) Consider $\mathscr{F}$ and $\mathscr{G}$ as in the statement of Proposition 3.4. Note that, in the first case, ${\rm T}\mathscr{F}={\rm T}\mathscr{G}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(\delta)$ is a subsheaf of $T{\mathbb{P}^{n}}$, hence $\delta\leq 1$, i.e., $c_{1}({\rm T}\mathscr{F})\leq c_{1}({\rm T}\mathscr{G})+1$. This is a simple constraint for having a subdistribution that splits.
2) If $\mathscr{F}$ is integrable, then the integrability of $\mathscr{G}$ is measured by the map ${\textstyle\bigwedge}^{2}{\rm T}\mathscr{G}\to{\rm T}\mathscr{F}/{\rm T}\mathscr{G}$ that sends $v\wedge w\mapsto[v,w]+{\rm T}\mathscr{G}$. Since ${\rm T}\mathscr{G}$ is locally free, this produces a section in $H^{0}({\textstyle\bigwedge}^{2}{\rm T}\mathscr{G}^{\vee}(\delta)\otimes\mathscr{I}_{W})$, with $W$ possibly empty. Then, $\mathscr{G}$ is automatically integrable if this cohomology vanishes.
Combining Proposition 3.2 and Proposition 3.4, one gets the following result.
Corollary 3.5.
Let $\mathscr{F}$ be a degree-$d$ foliation of dimension $k$ on $\mathbb{P}^{n}$}. Suppose that $h^{0}({\rm T}\mathscr{F})=k-1$ and $\mu_{\rm max}({\rm T}\mathscr{F})\leq 0$, then either ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus k-1}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(k-d)$ or $\mathfrak{fix}(\mathscr{F})$ defines a subfoliation of dimension k1k-1 and with singularities in codimension two.
In the next two subsections, we investigate foliations of dimensions 2 and 3 such that $0<h^{0}({\rm T}\mathscr{F})\leq\dim\mathscr{F}-1$.
3.1. 2-dimensional foliations tangent to a global vector field
For 2-dimensional foliations, $\rm T$$\mathscr{F}$ is a rank-2 sheaf. Then, Corollary 3.5 applies to any foliation of degree $d\geq 2$ such that $h^{0}({\rm T}\mathscr{F}(-1))=0$.
Remark 3.2.
A result by da Costa, see [13, Theorem C], says that foliations on $\mathbb{P}^{3}$} with locally-free tangent sheaf having a global section must split as a direct sum of line bundles. Using Corollary 3.5 we can extend this result to 22-dimensional foliations on $\mathbb{P}^{n}$}, $n\geq 3$.
Proposition 3.6.
Let $\mathscr{F}$ be a $2$-dimensional foliation of degree $d$ on $\mathbb{P}^{n}$} such that $h^{0}({\rm T}\mathscr{F})>0$. Then, either ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}(a)\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(1+a-d)$, $a\in\{0,1\}$, or $\rm T$$\mathscr{F}$ is not locally free.
Proof.
If $h^{0}({\rm T}\mathscr{F}(-1))>0$ then $\mathscr{F}$ is a linear pullback and ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}(1)\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(1-d)$. Suppose that $h^{0}({\rm T}\mathscr{F}(-1))=0$ and ${\rm T}\mathscr{F}\not\cong{\mathcal{O}}_{\mathbb{P}^{n}}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(2-d)$. Since $\operatorname{rk}{\rm T}\mathscr{F}=2$, it is enough to prove that $c_{3}({\rm T}\mathscr{F})\neq 0$. Consider the exact sequence from the first column of (3.1):
$0\longrightarrow{\mathcal{O}}_{\mathbb{P}^{n}}\longrightarrow{\rm T}\mathscr{F}\longrightarrow\mathscr{I}_{W}(2-d)\longrightarrow 0$
By Corollary 3.5 $W$ has pure codimension $2$. Furthermore, if $n>3$, then $W\cong{\mathbb{P}^{n-2}}$ is a linear subspace; and if $n=3$, then $W$ is either one line or a (possibly degenerate) disjoint union of two lines, see [18, §2.3] and Lemma 3.7 below.
If $n\geq 4$, we take a general $L\cong{\mathbb{P}^{3}}$ linearly embedded in $\mathbb{P}^{n}$} transverse to $W$ then the above sequence restricts to $L$ as
$0\longrightarrow\mathcal{O}_{L}\longrightarrow{\rm T}\mathscr{F}|_{L}\longrightarrow\mathscr{I}_{W\cap L}(2-d)\longrightarrow 0,$
$W\cap L$ is a line, and $c_{i}({\rm T}\mathscr{F}|_{L})=c_{i}({\rm T}\mathscr{F})$ for $i=1,2,3$. Therefore, we may assume that $n=3$.
By [21, Theorem 4.1], $c_{1}({\rm T}\mathscr{F})=2-d$, $c_{2}({\rm T}\mathscr{F})=\deg(W)$ and $c_{3}({\rm T}\mathscr{F})=2p_{a}(W)-2+\deg(W)(d+2)$. If $W$ is a line then $\deg(W)=1$ and $p_{a}(W)=0$, hence $c_{3}({\rm T}\mathscr{F})=d\geq 1$. If $W$ is (possibly the limit of) a disjoint union of two lines, $\deg(W)=2$ and $p_{a}(W)=-1$, hence $c_{3}({\rm T}\mathscr{F})=2d\geq 2$. ∎
We classify the possible subfoliations using Jordan normal forms.
Lemma 3.7.
Let $v\in H^{0}(T{\mathbb{P}^{n}})$ such that $v$ has zeros in codimension two. Then, up to a linear change of coordinates, either
  1. 1)
    $v=x_{0}\frac{\partial}{\partial x_{0}}+\lambda x_{1}\frac{\partial}{\partial x_{1}}$ with $\lambda\in\mathbb{C}^{*}$;
  2. 2)
    $v=(x_{0}+x_{1})\frac{\partial}{\partial x_{0}}+x_{1}\frac{\partial}{\partial x_{1}}$;
  3. 3)
    $v=x_{1}\frac{\partial}{\partial x_{0}}+x_{3}\frac{\partial}{\partial x_{2}}$;
  4. 4)
    $v=x_{1}\frac{\partial}{\partial x_{0}}+x_{2}\frac{\partial}{\partial x_{1}}$;
  5. 5)
    $v=x_{1}\frac{\partial}{\partial x_{0}}+x_{2}\frac{\partial}{\partial x_{2}}$.
And this allows us to conclude that these subfoliations are algebraically integrable.
Proposition 3.8.
Let $\mathscr{F}$ be a foliation of dimension $2$ and degree $d\geq 1$. Suppose that $h^{0}({\rm T}\mathscr{F}(-1))=0$ and $h^{0}({\rm T}\mathscr{F})>0$ and ${\rm T}\mathscr{F}\not\cong{\mathcal{O}}_{\mathbb{P}^{n}}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(2-d)$. For any $s\in H^{0}({\rm T}\mathscr{F})\setminus\{0\}$, the induced foliation by curves is algebraically integrable with singularities in codimension two.
Proof.
Let $\mathscr{G}$ be the foliation induced by a nonzero section $s\in H^{0}({\rm T}\mathscr{F})$. By Corollary 3.5, we know that $\mathscr{G}$ has singularities in codimension two. Suppose, aiming at a contradiction, that the general leaf of $\mathscr{G}$ is not algebraic. Then, due to Lemma 3.7, $\mathscr{G}$ is defined by either $v=x_{0}\frac{\partial}{\partial x_{0}}+\lambda x_{1}\frac{\partial}{\partial x_{1}}$ with $\lambda\not\in\mathbb{Q}$, or $v=(x_{0}+x_{1})\frac{\partial}{\partial x_{0}}+x_{1}\frac{\partial}{\partial x_{1}}$, or $v=x_{1}\frac{\partial}{\partial x_{0}}+x_{2}\frac{\partial}{\partial x_{2}}$. Let $\omega\in H^{0}(\Omega_{\mathbb{P}^{n}}^{n-2}(d+n-1))$ be a (n2)(n-2)-form defining $\mathscr{F}$.
Case 1: Suppose $v=x_{0}\frac{\partial}{\partial x_{0}}+\lambda x_{1}\frac{\partial}{\partial x_{1}}$ with $\lambda\not\in\mathbb{Q}$. Then $x_{0}\frac{\partial}{\partial x_{0}},x_{1}\frac{\partial}{\partial x_{1}}\in\mathfrak{afix}(\mathscr{F})$:
$\mathcal{L}_{x_{0}\frac{\partial}{\partial x_{0}}}\omega=\lambda_{0}\omega\quad\text{and}\quad\mathcal{L}_{x_{1}\frac{\partial}{\partial x_{1}}}\omega=\lambda_{1}\omega.$
Then $\omega$ is homogeneous of degree $\lambda_{j}$ on the variable $x_{j}$, $j=0,1$. Since $H^{0}({\rm T}\mathscr{F}(-1))=0$, we have $\lambda_{0}=\lambda_{1}=1$. Using that $\iota_{v}\omega=0$, one concludes that
$\omega=(x_{0}dx_{1}-\lambda x_{1}dx_{0})\wedge\alpha+x_{0}x_{1}\beta+dx_{0}\wedge dx_{1}\wedge\gamma$
where $\alpha$, $\beta$, and $\gamma$ do not depend on $x_{0},x_{1},dx_{0},dx_{1}$. From $\iota_{\operatorname{rad}}\omega=0$, one gets $\gamma=0$ and $\alpha=\frac{1}{\lambda-1}\iota_{\operatorname{rad}}\beta$. Thus, up to a constant,
$\omega=(x_{0}dx_{1}-\lambda x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\beta+(\lambda-1)x_{0}x_{1}\beta.$
Note that $\beta$ is a (n2)(n-2)-form in n1n-1 variables. Therefore, there exists a polynomial vector field $u=\sum_{i=2}^{n}g_{i}\frac{\partial}{\partial x_{i}}$ such that $g_{i}=g_{1}(x_{2},\dots,x_{n})$ is homogeneous of degree d1d-1 and
$\beta=\iota_{u}dx_{2}\wedge\dots\wedge dx_{n}.$
It follows that $\iota_{u}\omega=0$ and $[u,v]=0$. Thus ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(2-d)$ spanned by $v$ and $u$, which is a contradiction.
Case 2: If $v=(x_{0}+x_{1})\frac{\partial}{\partial x_{0}}+x_{1}\frac{\partial}{\partial x_{1}}$, we have that $x_{0}\frac{\partial}{\partial x_{0}}+x_{1}\frac{\partial}{\partial x_{1}},x_{1}\frac{\partial}{\partial x_{0}}\in\mathfrak{afix}(\mathscr{F})$. Then $\omega$ is homogeneous with respect to $\{x_{0},x_{1}\}$ and $\mathcal{L}_{x_{1}\frac{\partial}{\partial x_{0}}}\omega=0$, which, combined with $\iota_{\operatorname{rad}}\omega=0$, imply that
$\omega=(x_{1}dx_{0}-x_{0}dx_{1})\wedge\eta+dx_{0}\wedge dx_{1}\wedge\iota_{\operatorname{rad}}\eta+\gamma$
where $\eta$ and $\gamma$ do not depend on $x_{0},dx_{0},dx_{1}$. By homogeneity, $\eta$ does not depend on $x_{1}$ and $\gamma=x_{1}^{2}\gamma^{\prime}$. Then
$0=\iota_{v}\omega=x_{1}^{2}\eta+((x_{0}+x_{1})dx_{1}-x_{1}dx_{0})\wedge\iota_{v}\iota_{\operatorname{rad}}\eta,$
and we conclude that $\eta=0$. Therefore, $\omega$ does not depend on $x_{0},x_{1},dx_{0},dx_{1}$. This implies $h^{0}({\rm T}\mathscr{F}(-1))>0$ which is a contradiction.
Case 3: If $v=x_{1}\frac{\partial}{\partial x_{0}}+x_{2}\frac{\partial}{\partial x_{2}}$, we have that $x_{2}\frac{\partial}{\partial x_{2}},x_{1}\frac{\partial}{\partial x_{0}}\in\mathfrak{afix}(\mathscr{F})$. Then, again,
$\omega=(x_{1}dx_{0}-x_{0}dx_{1})\wedge\eta+dx_{0}\wedge dx_{1}\wedge\iota_{\operatorname{rad}}\eta+\gamma$
where $\eta$ and $\gamma$ do not depend on $x_{0},dx_{0},dx_{1}$. And $\omega$ is homogeneous with respect to $x_{2}$. Then
$\iota_{v}\omega$ $=x_{1}^{2}\eta-x_{2}(x_{1}dx_{0}-x_{0}dx_{1})\wedge\iota_{\frac{\partial}{\partial x_{2}}}\eta+x_{1}dx_{1}\wedge\iota_{\operatorname{rad}}\eta+x_{2}dx_{0}\wedge dx_{1}\wedge\iota_{\frac{\partial}{\partial x_{2}}}\iota_{\operatorname{rad}}\eta+x_{2}\iota_{\frac{\partial}{\partial x_{2}}}\gamma$
$=dx_{0}\wedge(-x_{1}x_{2}\iota_{\frac{\partial}{\partial x_{2}}}\eta)+dx_{1}\wedge(x_{0}x_{2}\iota_{\frac{\partial}{\partial x_{2}}}\eta+x_{1}\iota_{\operatorname{rad}}\eta)+dx_{0}\wedge dx_{1}\wedge(x_{2}\iota_{\frac{\partial}{\partial x_{2}}}\iota_{\operatorname{rad}}\eta)+$
$\quad+x_{1}^{2}\eta+x_{2}\iota_{\frac{\partial}{\partial x_{2}}}\gamma=0$
Therefore, $\iota_{\frac{\partial}{\partial x_{2}}}\eta=\iota_{\operatorname{rad}}\eta=x_{1}^{2}\eta+x_{2}\iota_{\frac{\partial}{\partial x_{2}}}\gamma=0$ and
$\omega=(x_{1}x_{2}dx_{0}-x_{0}x_{2}dx_{1}-x_{1}^{2}dx_{2})\wedge\theta+x_{2}\rho$
where $\eta=x_{2}\theta$ and $\gamma=-x_{1}^{2}dx_{2}\wedge\theta+x_{2}\rho$. Moreover, $\theta$ and $\rho$ do not depend on $x_{0},x_{2},dx_{0},dx_{1},dx_{2}$. Note that $\iota_{u}\omega=0$ for $u=x_{0}\frac{\partial}{\partial x_{0}}+x_{1}\frac{\partial}{\partial x_{1}}$, which commutes with $v$. Therefore, ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus 2}$, which is a contradiction. ∎
In light of the previous proposition, we need to understand foliations tangent to algebraic actions with codimension-two fixed points. We start with additive $\mathbb{C}$}-actions, which correspond to cases 3) and 4) in Lemma 3.7 up to a linear change of coordinates.
Proposition 3.9.
Let $\mathscr{F}$ be a foliation of dimension $2$ and degree $2$ on $\mathbb{P}^{n}$} tangent to a global $\mathbb{C}$}-action with codimension-two fixed points. Then $\mathscr{F}\in{\rm Log}(1^{n-2},2;2,n)$.
Proof.
We fix an action tangent to a vector field in either 3) or 4) in Lemma 3.7. Both cases follow the same argument, so we only prove one. In case 3), $v=x_{1}\frac{\partial}{\partial x_{0}}+x_{3}\frac{\partial}{\partial x_{2}}$. Then it has a rational first integral $\rho\colon{\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}}(1^{n-2},2)$ given by
$\rho(x_{0}:\dots:x_{n})=(x_{1}:x_{3}:x_{4}:\dots:x_{n}:x_{0}x_{3}-x_{1}x_{2}).$
Then $\mathscr{F}=\rho^{*}\mathscr{H}$ for $\mathscr{H}$ a foliation by curves on ${\mathbb{P}}(1^{n-2},2)$. Note that $\rho$ does not ramify in codimension one, hence
${\mathcal{O}}_{\mathbb{P}^{n}}(n+1)=\det({\rm N}\mathscr{F})=\rho^{*}\det({\rm N}\mathscr{H})={\mathcal{O}}_{\mathbb{P}^{n}}(n+\deg\mathscr{H}).$
Thus $\deg(\mathscr{H})=1$ and $\mathscr{H}$ is defined by a global vector field. As we observed in Remark 2.2, $\mathscr{F}\in{\rm Log}(1^{n-2},2;2,n)$. ∎
Actions of the multiplicative group $\mathbb{C}$$^{\ast}$ with codimension-two fixed locus are, up to a linear change of coordinates, induced by vector fields as in 1) in Lemma 3.7, $v=x_{0}\frac{\partial}{\partial x_{0}}+\mu x_{1}\frac{\partial}{\partial x_{1}}\in H^{0}(T{\mathbb{P}^{n}})$ with $\mu\in\mathbb{Q}$. The next result classifies the $2$-dimensional foliations invariant under such actions. Note that we don’t impose any restriction on $\mu\in{\mathbb{C}}^{*}$ a priori.
Proposition 3.10.
Let $\omega\in H^{0}(\Omega_{{\mathbb{P}^{n}}}^{n-2}(n+1))$ not vanishing in codimension one and let $v=x_{0}\frac{\partial}{\partial x_{0}}+\mu x_{1}\frac{\partial}{\partial x_{1}}\in H^{0}(T{\mathbb{P}^{n}})$ be such that $\iota_{v}\omega=0$ and $\iota_{v}d\omega=a\omega$ for $a\in{\mathbb{C}}$ a constant. Then $\omega$ defines a foliation $\mathscr{F}$ of one of the following types:
  1. 1)
    $a=\mu+1$, $\mu\neq-1$ and $\mathscr{F}\in{\rm Log}(1^{n+1};2,n)$;
  2. 2)
    $a=0$, $\mu=-1$ and $\mathscr{F}\in{\rm Log}(1^{n-1},2;2,n)$;
  3. 3)
    $a=0$, $\mu=-2$ and $\mathscr{F}\in{\rm Rat}(1^{n-2},3;n)$;
  4. 4)
    $\mu\in\{1,2\}$ and $a=\mu+2$, and $\mathscr{F}$ is induced by an action of $\mathfrak{aff}({\mathbb{C}})$.
Proof.
Since $\iota_{v}\omega=0$ we may write
$\omega=(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\beta+(\mu-1)x_{0}x_{1}\beta\,+\,\gamma,$
where $\beta$ and $\gamma$ do not depend on $dx_{0}$ nor $dx_{1}$ and $\iota_{\operatorname{rad}}\gamma=0$. Then $\beta$ and $\gamma$ are (n2)(n-2)-forms depending on n1n-1 basic differentials, which means that there exists a degree-two homogeneous polynomial $q$ and a linear vector field $u=\sum_{j=2}^{n}b_{j}\frac{\partial}{\partial x_{j}}$ such that $\beta=\iota_{u}\Theta$ and $\gamma=q\,\iota_{\operatorname{rad}}\Theta$, where $\Theta=dx_{2}\wedge\dots\wedge dx_{n}$. Hence,
$\omega=(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\iota_{u}\Theta+(\mu-1)x_{0}x_{1}\iota_{u}\Theta\,+\,q\,\iota_{\operatorname{rad}}\Theta.$
We might assume that $u$ is not proportional to the radial in the last n1n-1 variables, i.e., $u\wedge\sum_{j=2}^{n}x_{j}\frac{\partial}{\partial x_{j}}\neq 0$, otherwise $\omega$ would vanish in codimension one.
By hypothesis, there exists $a\in{\mathbb{C}}$ such that $\iota_{v}d\omega=a\omega$. Computing the LHS, we will see that there are few possibilities for $a$. Start with expanding
$d\omega$ $=(\mu+1)dx_{0}\wedge dx_{1}\wedge\iota_{\operatorname{rad}}\iota_{u}\Theta-(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge d\iota_{\operatorname{rad}}\iota_{u}\Theta+$
$\quad+(\mu-1)(x_{0}dx_{1}+x_{1}dx_{0})\wedge\iota_{u}\Theta+(\mu-1)x_{0}x_{1}d\iota_{u}\Theta+dq\wedge\iota_{\operatorname{rad}}\Theta+\,q\,d\iota_{\operatorname{rad}}\Theta$
then,
$\iota_{v}d\omega$ $=(\mu+1)(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\iota_{u}\Theta+(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge\iota_{v}d\iota_{\operatorname{rad}}\iota_{u}\Theta+$
$\quad+(\mu-1)(\mu+1)x_{0}x_{1}\iota_{u}\Theta+(\mu-1)x_{0}x_{1}\iota_{v}d\iota_{u}\Theta+v(q)\iota_{\operatorname{rad}}\Theta$
$=(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge((\mu+1)\iota_{\operatorname{rad}}\iota_{u}\Theta+\iota_{v}d\iota_{\operatorname{rad}}\iota_{u}\Theta)+$
$\quad+(\mu-1)x_{0}x_{1}((\mu+1)\iota_{u}\Theta+\iota_{v}d\iota_{u}\Theta)+v(q)\iota_{\operatorname{rad}}\Theta$
and by $\iota_{v}d\omega=a\omega$ imposing we get that
$\iota_{v}d\iota_{\operatorname{rad}}\iota_{u}\Theta$ $=(a-\mu-1)\iota_{\operatorname{rad}}\iota_{u}\Theta,$
$\iota_{v}d\iota_{u}\Theta$ $=(a-\mu-1)\iota_{u}\Theta+c\,\iota_{\operatorname{rad}}\Theta$
$v(q)$ $=a\,q-(\mu-1)c\,x_{0}x_{1}$
where $c\in{\mathbb{C}}$ and the second equation follows from the first.
Notice that $\iota_{\operatorname{rad}}\iota_{u}\Theta=\sum_{i,j}p_{ij}dx_{2}\wedge\dots\wedge\widehat{dx_{i}}\wedge\widehat{dx_{j}}\wedge\dots\wedge dx_{n}$, where each $p_{ij}$ is a quadratic polynomial not having monomials $x_{0}^{2}$, $x_{0}x_{1}$ nor $x_{1}^{2}$. Moreover,
$\iota_{v}d\iota_{\operatorname{rad}}\iota_{u}\Theta=\sum_{i,j}v(p_{ij})dx_{2}\wedge\dots\wedge\widehat{dx_{i}}\wedge\widehat{dx_{j}}\wedge\dots\wedge dx_{n}$
and we only need to compute the action of $v$ on degree two polynomials. From $v(x_{0})=x_{0}$, $v(x_{1})=\mu x_{1}$ and $v(x_{j})=0$ for $j\geq 2$ we get that
(3.2) $\frac{v(x_{i}x_{j})}{x_{i}x_{j}}=\begin{cases}2,&i=j=0\\ 2\mu,&i=j=1\\ (\mu+1),&i=0,\,j=1\\ 1,&i=0,\,j\geq 2\\ \mu,&i=1,\,j\geq 2\\ 0,&i,j\geq 2\end{cases}$
Then, to have $\iota_{v}d\iota_{\operatorname{rad}}\iota_{u}\Theta=(a-\mu-1)\iota_{\operatorname{rad}}\iota_{u}\Theta$, we need $a-\mu-1\in\{0,1,\mu\}$.
Case 1.0:
Suppose that $a-\mu-1=0$ and that $\mu\not\in\{-1,1\}$. Then $\iota_{v}d\iota_{u}\Theta=c\,\iota_{\operatorname{rad}}\Theta$ implies $c=0$. Thus $v(q)=(\mu+1)q$. From (3.2) and $\mu\not\in\{-1,1\}$, $q=c^{\prime}x_{0}x_{1}$ for $c^{\prime}\in{\mathbb{C}}$ a constant. Then
$\omega=(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\iota_{u}\Theta+x_{0}x_{1}((\mu-1)\iota_{u}\Theta\,+c^{\prime}\,\iota_{\operatorname{rad}}\Theta).$
Consider $w=c^{\prime}\,x_{1}\frac{\partial}{\partial x_{1}}+u$ and notice that
$\iota_{w}\iota_{v}\iota_{\operatorname{rad}}(dx_{0}\wedge\dots\wedge dx_{n})=\iota_{w}\iota_{v}\iota_{\operatorname{rad}}(dx_{0}\wedge dx_{1}\wedge\Theta)=\omega.$
Also, since $u$ does not depend on $x_{0}$ nor $x_{1}$, $[v,w]=0$. Thus $\omega$ defines a foliation $\mathscr{F}$ on $\mathbb{P}^{n}$} with trivial tangent sheaf ${\rm T}\mathscr{F}={\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus 2}$.
Case 1.1:
Suppose that $a-\mu-1=0$ and that $\mu=1$. As in the previous case, $c=0$ and $u$ does not depend on $x_{0}$ nor $x_{1}$. But now $v(q)=2q$ implies that $q=q_{0}x_{0}^{2}+2q_{1}x_{0}x_{1}+q_{2}x_{1}^{2}$. Then
$\omega=(x_{0}dx_{1}-x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\iota_{u}\Theta+(q_{0}x_{0}^{2}+2q_{1}x_{0}x_{1}+q_{2}x_{1}^{2})\iota_{\operatorname{rad}}\Theta.$
Consider the vector field
$w=\left(-q_{1}x_{0}-q_{2}x_{1}\right)\frac{\partial}{\partial x_{0}}+\left(q_{0}x_{0}+q_{1}x_{1}\right)\frac{\partial}{\partial x_{1}}+u.$
As in the previous case, we have that $\omega=\iota_{w}\iota_{v}\iota_{\operatorname{rad}}(dx_{0}\wedge dx_{1}\wedge\Theta)$, and it defines a foliation whose tangent sheaf is trivial.
Case 1.2:
Suppose that $a-\mu-1=0$ and that $\mu=-1$. Again, we have $c=0$ but now $v(q)=0$, which implies that $q=q_{0}x_{0}x_{1}+q^{\prime}(x_{2},\dots,x_{n})$. Then
$\omega=d(x_{0}x_{1})\wedge\iota_{\operatorname{rad}}\iota_{u}\Theta-2x_{0}x_{1}\iota_{u}\Theta\,+\,q\,\iota_{\operatorname{rad}}\Theta.$
Then $\omega$ defines a foliation on $\mathbb{P}^{n}$} which is the pullback of the foliation on ${\mathbb{P}}(2,1^{(n-1)})$ given by the vector field
$m=(q_{0}y_{1}+q^{\prime}(y_{2},\dots,y_{n}))\frac{\partial}{\partial y_{1}}+\sum_{j=2}^{n}b_{j}(y)\frac{\partial}{\partial y_{j}}$
via the map defined by $y_{1}=x_{0}x_{1}$ and $y_{j}=x_{j}$ for $j\geq 2$; recall that $u=\sum_{j=2}^{n}b_{j}(x)\frac{\partial}{\partial x_{j}}$.
Case 2.0:
Suppose that $a-\mu-1=1$ and that $\mu\not\in\{-2,-1,2\}$. Then $\iota_{v}\iota_{\operatorname{rad}}\iota_{u}\Theta=\iota_{\operatorname{rad}}\iota_{u}\Theta$ and $x_{0}$ must divide $\iota_{\operatorname{rad}}\iota_{u}\Theta$. Hence, there exist $\xi$ and $c^{\prime}$ such that $\iota_{u}\Theta=x_{0}\xi+c^{\prime}\iota_{\operatorname{rad}}\Theta$. Note that $c^{\prime}=-c$. Then
$\omega=x_{0}(x_{0}dx_{1}-\mu x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\xi+(\mu-1)x_{0}^{2}x_{1}\xi+(-c(\mu-1)x_{0}x_{1}+q)\iota_{\operatorname{rad}}\Theta.$
On the other hand, $v(q)=(\mu+2)q-(\mu-1)c\,x_{0}x_{1}$ implies that $q=c(\mu-1)\,x_{0}x_{1}$. Hence, $\omega$ vanishes along $V(x_{0})$ which contradicts our hypothesis.
Case 2.1:
Suppose that $a-\mu-1=1$ and that $\mu=-2$. Then, $v(q)=3cx_{0}x_{1}$ which implies that $q=-3cx_{0}x_{1}+q^{\prime}(x_{2},\dots,x_{n})$. Hence,
$\omega$ $=x_{0}(x_{0}dx_{1}+2x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\xi-3x_{0}^{2}x_{1}\xi+(3c\,x_{0}x_{1}+q)\iota_{\operatorname{rad}}\Theta=$
$=x_{0}[(x_{0}dx_{1}+2x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\xi-3x_{0}x_{1}\xi]+q^{\prime}\iota_{\operatorname{rad}}\Theta.$
Note that $\xi$ is a constant (n2)(n-2)-from on $\{dx_{2},\dots,dx_{n}\}$. Then, up to a linear change on the last n1n-1 coordinates, we may assume that $\xi=dx_{3}\wedge\dots\wedge dx_{n}$. Then
$\omega=\iota_{\operatorname{rad}}((q^{\prime}dx_{2}-d(x_{0}^{2}x_{1}))\wedge dx_{3}\wedge\dots\wedge dx_{n}).$
Now consider $f(x_{2},\dots,x_{n})$ a degree $3$ polynomial such that $\frac{\partial f}{\partial x_{2}}=q^{\prime}$. It follows that
$\omega=\iota_{\operatorname{rad}}(d(f-x_{0}^{2}x_{1})\wedge dx_{3}\wedge\dots\wedge dx_{n}),$
and $\omega$ defines a rational foliation of type $(3,1^{(n-2)})$.
Case 2.2:
Suppose that $a-\mu-1=1$ and that $\mu=2$. Then, $v(q)=4q-cx_{0}x_{1}$ which implies $q=q_{0}x_{1}^{2}+c\,x_{0}x_{1}$. Then
$\omega$ $=x_{0}(x_{0}dx_{1}-2x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\xi+x_{0}^{2}x_{1}\xi+(-cx_{0}x_{1}+q)\iota_{\operatorname{rad}}\Theta$
$=x_{0}(x_{0}dx_{1}-2x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\xi+x_{0}^{2}x_{1}\xi+q_{0}x_{1}^{2}\iota_{\operatorname{rad}}\Theta$
As in the previous case, we may assume $\xi=dx_{3}\wedge\dots\wedge dx_{n}$, hence
$\omega$ $=x_{0}(x_{0}dx_{1}-2x_{1}dx_{0})\wedge\iota_{\operatorname{rad}}\xi+x_{0}^{2}x_{1}\xi+q_{0}x_{1}^{2}(x_{2}\xi-dx_{2}\wedge\iota_{\operatorname{rad}}\xi)$
$=(x_{0}(x_{0}dx_{1}-2x_{1}dx_{0})-q_{0}x_{1}^{2}dx_{2})\wedge\iota_{\operatorname{rad}}\xi+(x_{0}^{2}x_{1}+q_{0}x_{1}^{2}x_{2})\xi$
$=\iota_{\operatorname{rad}}((x_{0}^{2}dx_{1}-2x_{1}x_{0}dx_{0}-q_{0}x_{1}^{2}dx_{2})\wedge dx_{3}\wedge\dots\wedge dx_{n}).$
Note that $q_{0}\neq 0$ unless $\omega$ vanishes along $V(x_{0})$. Define the vector field $w=\frac{q_{0}}{2}x_{1}\frac{\partial}{\partial x_{0}}+x_{0}\frac{\partial}{\partial x_{2}}$. It follows that
$\omega=\iota_{\operatorname{rad}}\iota_{w}\iota_{v}((x_{0}^{2}dx_{1}-2x_{1}x_{0}dx_{0}-q_{0}x_{1}^{2}dx_{2})\wedge dx_{3}\wedge\dots\wedge dx_{n}).$
Also notice that $[v,w]=w$ hence $\omega$ defines a foliation associated with an algebraic action of the affine Lie algebra.
Case 2.3:
Suppose that $a-\mu-1=1$ and that $\mu=-1$. Again $x_{0}$ divides $\iota_{\operatorname{rad}}\iota_{u}\Theta$. But now $v(q)=q+2c\,x_{0}x_{1}$, which implies that $x_{0}$ divides $q$. Then $\omega$ vanishes along $V(x_{0})$ which is absurd.
Case 2.4:
Suppose that $a-\mu-1=1$ and that $\mu=1$. Then $v(q)=3q$ implies $q=0$. Then,
$\omega=\iota_{\operatorname{rad}}\iota_{v}\iota_{u}(dx_{0}\wedge\dots\wedge dx_{n}).$
On the other hand, $\iota_{v}d\iota_{\operatorname{rad}}\iota_{u}\Theta=\iota_{\operatorname{rad}}\iota_{u}\Theta$ implies that $u=u^{\prime}+c^{\prime}\sum_{j=2}x_{j}\frac{\partial}{\partial x_{j}}$, where the coefficients of $u^{\prime}$ depend uniquely on $x_{0}$ and $x_{1}$. Note that we can replace $u$ with $u^{\prime}$ in the definition of $\omega$. Moreover, $v,u^{\prime}]=u^{\prime}$ and $\omega$ defines a foliation tangent to an action of $\mathfrak{aff}({\mathbb{C}})$.
Case 3:
Suppose that $a-\mu-1=\mu$. We may exchange $x_{0}$ and $x_{1}$ so that we get new parameters $(\mu^{\prime},a^{\prime})=(1/\mu,a/\mu)$. For $a=2\mu+1$ we get $a^{\prime}=\mu^{\prime}+2$. Then we reduce to one of the previous cases. ∎
3.2. 3-dimensional foliations with nonstable tangent sheaf.
We focus on the case $c_{1}({\rm T}\mathscr{F})=0$, i.e., degree three. Note that, in this case, $\rm T$$\mathscr{F}$ is stable if and only if $h^{0}({\rm T}\mathscr{F})=h^{0}({\rm T}\mathscr{F}^{\vee})=0$.
Proposition 3.11.
Let $\mathscr{F}$ be a 3-dimensional foliation on $\mathbb{P}^{n}$} of degree $3$. Suppose that $\rm T$$\mathscr{F}$ is not semistable, then either
  • $\mathscr{F}$ is a linear pullback of a degree-3 foliation on $\mathbb{P}^{n-1}$} or $\mathbb{P}^{n-2}$}, or
  • $\mathscr{F}$ is a pullback of a degree-3 foliation on ${\mathbb{P}}(1^{n-2},2)$.
Proof.
Since $\rm T$$\mathscr{F}$ is not semistable, either $h^{0}({\rm T}\mathscr{F}(-1))\neq 0$ or there exists $G\subset{\rm T}\mathscr{F}$ such that $c_{1}(G)>0$. The latter is equivalent to $h^{0}({\rm T}\mathscr{F}^{\vee}(-1))\neq 0$. In the first case, $\mathscr{F}$ is a linear pullback of a degree-3 foliation on $\mathbb{P}^{n-h}$} for $h=h^{0}({\rm T}\mathscr{F}(-1))\leq 2$. This gives our first alternative.
Assume $h^{0}({\rm T}\mathscr{F}(-1))=0$ and $h^{0}({\rm T}\mathscr{F}^{\vee}(-1))\neq 0$. This corresponds to an exact sequence
$0\longrightarrow{\rm T}\mathscr{G}\longrightarrow{\rm T}\mathscr{F}\longrightarrow\mathscr{I}_{W}(-1)\longrightarrow 0$
for some subdistribution $\mathscr{G}$ and some subscheme $W$. Note that $c_{1}({\rm T}\mathscr{G})=1$ and $h^{0}({\rm T}\mathscr{G}(-1))=0$. Then, ${\rm T}\mathscr{G}$ is stable and a maximal destabilizing subsheaf for $\rm T$$\mathscr{F}$. Hence, $\mathscr{G}\subset\mathscr{F}$ is a subfoliation. By the classification of degree-one foliations [26, Theorem 6.2], $\mathscr{G}$ is defined by a rational map ${\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}}(1^{n-2},2)$ given by n2n-2 linear polynomials and one quadratic. This proves the second alternative. ∎
Proposition 3.12.
Let $\mathscr{F}$ be a 3-dimensional foliation on $\mathbb{P}^{n}$} of degree $3$. Suppose $\rm T$$\mathscr{F}$ is strictly semistable and indecomposable, i.e., not a direct sum of two other sheaves. Then either
  1. 1)
    $\mathscr{F}$ is tangent to a global vector field singular in codimension at most 3;
  2. 2)
    $\mathscr{F}$ is tangent to a 2-dimensional Lie algebra action singular in codimension 2;
  3. 3)
    $\mathscr{F}$ admits a 2-dimensional subfoliation $\mathscr{G}$ of degree 2 such that ${\rm T}\mathscr{G}$ is stable;
  4. 4)
    $\rm T$$\mathscr{F}$ is defined by a nontrivial extension $0\to{\mathcal{O}}_{\mathbb{P}^{n}}\to{\rm T}\mathscr{F}^{\vee}\to Q$, where $Q$ is a stable reflexive sheaf.
Proof.
First, $\rm T$$\mathscr{F}$ being strictly semistable means that $h^{0}({\rm T}\mathscr{F}(-1))=h^{0}({\rm T}\mathscr{F}^{\vee}(-1))=0$ and either $h^{0}({\rm T}\mathscr{F})\neq 0$ or $h^{0}({\rm T}\mathscr{F}^{\vee})\neq 0$. By Lemma 3.1, $h^{0}({\rm T}\mathscr{F})\leq 2$; otherwise ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus 3}$. If $h^{0}({\rm T}\mathscr{F})=1$, then we have a global section $\sigma$ and the corresponding vector field $v_{\sigma}$. If $v_{\sigma}$ vanishes in codimension at least $4$, then ${\rm T}\mathscr{F}={\mathcal{O}}_{\mathbb{P}^{n}}\oplus{\rm T}\mathscr{G}$, by de Rham-Saito division, see [25, §1]. Hence, $v_{\sigma}$ must vanish in codimension at most $3$. This gives the first alternative.
Suppose $h^{0}({\rm T}\mathscr{F})=2$. By Proposition 3.2, $\mathscr{F}$ admits a subfoliation $\mathscr{G}$ such that ${\rm T}\mathscr{G}={\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus 2}$. By Proposition 3.4, ${\rm T}\mathscr{F}/{\rm T}\mathscr{F}=\mathscr{I}_{W}$ for $\emptyset\neq W\subset\operatorname{Sing}(\mathscr{G})$ of codimension two. This gives the second alternative.
Next, suppose $h^{0}({\rm T}\mathscr{F})=0$. Then $h^{0}({\rm T}\mathscr{F}^{\vee})=1$. Indeed, if $h^{0}({\rm T}\mathscr{F}^{\vee})\geq 3$ then ${\rm T}\mathscr{F}={\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus 3}$, and if $h^{0}({\rm T}\mathscr{F}^{\vee})=2$ then $h^{0}({\rm T}\mathscr{F})\geq 1$. Let $\sigma\in H^{0}({\rm T}\mathscr{F}^{\vee})$. We get the following exact sequence.
(3.3) $0\longrightarrow{\rm T}\mathscr{G}\longrightarrow{\rm T}\mathscr{F}\stackrel{{\scriptstyle\sigma^{\vee}}}{{\longrightarrow}}\mathscr{I}_{W}\longrightarrow 0$
where ${\rm T}\mathscr{G}$ is stable of rank two, and $W$ is the degeneracy locus of $\sigma$. If $\mathscr{G}$ is integrable, we get the third alternative. Otherwise, the Lie bracket induces a nonzero morphism ${\textstyle\bigwedge}^{2}{\rm T}\mathscr{G}\to\mathscr{I}_{W}$. Note that this morphism kills the torsion part $T$ of ${\textstyle\bigwedge}^{2}{\rm T}\mathscr{G}$ and ${\textstyle\bigwedge}^{2}{\rm T}\mathscr{G}/T=\mathscr{I}_{Y}$ for $Y$ supported at the singular locus of ${\rm T}\mathscr{G}$. Taking double duals, one sees that any morphism $\mathscr{I}_{Y}\to\mathscr{I}_{W}$ is given by multiplication with a constant. We conclude that $W\subset Y$. In particular, $\operatorname{codim}W\geq 3$. Recall that ${\mathcal{E}}{\it xt}^{1}(\mathscr{I}_{W},{\mathcal{O}}_{\mathbb{P}^{n}})$ is supported on the codimension-two part of $W$, which is empty. Therefore, dualizing (3.3) one gets
$0\longrightarrow{\mathcal{O}}_{\mathbb{P}^{n}}\longrightarrow{\rm T}\mathscr{F}^{\vee}\longrightarrow{\rm T}\mathscr{G}^{\vee}\longrightarrow{\mathcal{E}}{\it xt}^{1}(\mathscr{I}_{W},{\mathcal{O}}_{\mathbb{P}^{n}})=0$
proving the last alternative. ∎
By a result of Druel, a foliation with stable tangent sheaf $c_{1}=0$ and dimension $\leq 3$ must be algebraically integrable. Thus, in the third alternative, the foliation $\mathscr{F}$ is a pullback of a foliation by curves.
An interesting question is whether an analog of Proposition 3.6 holds for 3-dimensional foliations such that $h^{0}({\rm T}\mathscr{F})\geq 2$. Following the same strategy, one is led to compute $\mathscr{G}$, the subfoliation defined by $\mathfrak{fix}(\mathscr{F})$. The computations of $\mathfrak{afix}(\mathscr{F})$ become too cumbersome, and we didn’t pursue this investigation.
4. Main result
In this section, we prove Theorem A. We start with the case of two-dimensional foliations.
Theorem 4.1.
Let $\mathscr{F}$ be a two-dimensional foliation on $\mathbb{P}^{n}$} of degree two. Then one of the following holds:
  1. 1)
    $\rm T$$\mathscr{F}$ is stable and $\mathscr{F}$ is algebraically integrable with uniruled leaves;
  2. 2)
    $\rm T$$\mathscr{F}$ is strictly semistable and $\mathscr{F}$ belongs to either
    • an Affine-Lie component ${\rm ALie}(d_{1},\dots,d_{r},2;n)$;
    • one of the rational components ${\rm Rat}(1^{n-2},2,2;n)$ or ${\rm Rat}(1^{n-1},3;n)$; or
    • one of the logarithmic components ${\rm Log}(1^{n-1},2;2,n)$ or ${\rm Log}(1^{n+1};2,n)$.
  3. 3)
    $\rm T$$\mathscr{F}$ is not semistable and $\mathscr{F}$ is a linear pullback of a degree-$2$ foliation by curves on $\mathbb{P}^{n-1}$}.
Proof.
Given a foliation $\mathscr{F}$ as in the statement of the theorem, its tangent sheaf $\rm T$$\mathscr{F}$ is reflexive of rank two with zero first Chern class. Then $\rm T$$\mathscr{F}$ is semistable if and only if $h^{0}({\rm T}\mathscr{F}(-1))=0$, and stable if and only if $h^{0}({\rm T}\mathscr{F})=0$. If $\rm T$$\mathscr{F}$ is not semistable, then ${\rm T}\mathscr{F}\cong{\mathcal{O}}_{\mathbb{P}^{n}}(1)\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(-1)$ and $\mathscr{F}$ is a linear pullback of a degree-$2$ foliation by curves on $\mathbb{P}^{n-1}$}. This proves item 3).
If $\rm T$$\mathscr{F}$ is stable then $c_{2}({\rm T}\mathscr{F})>0$, otherwise it would be trivial, see [20, Lemma 2.5]. Since $\mathbb{P}^{n}$} is smooth and simply connected, it follows from Druel’s results [16, Theorem 6.1 and Proposition 8.4] that $\mathscr{F}$ is algebraically integrable. Due to [27, Theorem 5.13], $\mathscr{F}$ has uniruled leaves. This proves the first item.
It remains to understand the case where $\rm T$$\mathscr{F}$ is strictly semistable. If ${\rm T}\mathscr{F}={\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus 2}$ then $\mathscr{F}$ is associated with a Lie algebra action, and thus it falls in either ${\rm Log}(1^{n+1};2,n)$ or an Affine-Lie component from Theorem 2.3. If ${\rm T}\mathscr{F}\not\cong{\mathcal{O}}_{\mathbb{P}^{n}}^{\oplus 2}$ then Proposition 3.8 says that $\mathscr{F}$ is tangent to a one-dimensional algebraic action with codimension-two fixed locus.
If $\mathscr{F}$ is tangent to an additive action, due to Proposition 3.9, then $\mathscr{F}\in{\rm Log}(1^{n-2},2;2,n)$. If $\mathscr{F}$ is tangent to a multiplicative action, Proposition 3.10 implies that $\mathscr{F}$ must fit in one of the cases already listed. ∎
Next, we treat the case of dimension $k>2$. Then, the positivity of $\rm T$$\mathscr{F}$ simplifies the classification. Indeed, $c_{1}({\rm T}\mathscr{F})=k-2$, and $\mathscr{F}$ is a Mukai foliation, see [2]. We have the following theorem (cf. Theorem A).
Theorem 4.2.
Let $\mathscr{F}$ be a degree-$2$ foliation on $\mathbb{P}^{n}$} of dimension $k>2$. Then one of the following holds:
  1. 1)
    $\mathscr{F}$ is algebraically integrable;
  2. 2)
    $\mathscr{F}\in{\rm Log}(1^{n-k+1},2;k,n)$ is a logarithmic foliation of type $(1^{n-r+1},2)$;
  3. 3)
    $\mathscr{F}\in{\rm Log}(1^{n+1};k,n)$ is a logarithmic foliation of type $(1^{n+1})$;
  4. 4)
    $\mathscr{F}\in{\rm ALie}(r_{1},\dots,r_{m},k;n)$ is an Affine-Lie foliation; or
  5. 5)
    $\mathscr{F}\in{\rm LPB}(2,k,n)$ is a linear pullback of a foliation by curves in $\mathbb{P}^{n-k+1}$}.
The proof of this result will follow from Proposition 4.3 below. We begin assuming that $\mathscr{F}$ is not a linear pullback of a foliation on $\mathbb{P}^{r}$} with $r<n$. This is equivalent to assuming that $\mathscr{F}$ has no subfoliation of degree 0 or that $H^{0}({\rm T}\mathscr{F}(-1))=0$. Indeed, for any foliation $\mathscr{G}$ of degree 0, we have ${\rm T}\mathscr{G}\cong{\mathcal{O}}_{\mathbb{P}^{n}}(1)^{\oplus\dim\mathscr{G}}$.
Proposition 4.3.
Let $\mathscr{F}$ a degree-$2$ foliation on $\mathbb{P}^{n}$} of dimension $k>2$. If $H^{0}({\rm T}\mathscr{F}(-1))=0$ then either
  • $\mathscr{F}$ is algebraically integrable, or
  • $\mathscr{F}\in{\rm Log}(1^{n-k+1,2};k,n)$ is a logarithmic foliation of type $(1^{n-r+1},2)$.
Proof.
Suppose that $\mathscr{F}$ is not algebraically integrable. By [1, Proposition 7.5] there exists a subfoliation $\mathscr{G}\subset\mathscr{F}$ such that $c_{1}({\rm T}\mathscr{G})\geq c_{1}({\rm T}\mathscr{F})$, hence,
$\dim\mathscr{G}-\deg\mathscr{G}\geq k-2>\dim\mathscr{G}-2$
which implies that $\deg\mathscr{G}\leq 1$. Since $H^{0}({\rm T}\mathscr{F}(-1))=0$, we have $\deg\mathscr{G}=1$ and $H^{0}({\rm T}\mathscr{G}(-1))=0$; hence, $\dim\mathscr{G}=k-1$. By Theorem 2.1, $\mathscr{G}$ is defined by a rational map $\rho\colon{\mathbb{P}^{n}}\dashrightarrow{\mathbb{P}}(1^{n-k+1},2)$ given by nk+1n-k+1 linear forms and one quadratic form. It follows that $\mathscr{F}=\rho^{*}\mathscr{H}$ for some (purely transcendental) foliation by curves $\mathscr{H}$ on ${\mathbb{P}}(1^{n-k+1},2)$. Since $\rho$ does not ramify in codimension one, $\det({\rm N}\mathscr{F})=\rho^{*}\det({\rm N}\mathscr{H})$. Hence, $\mathscr{H}$ is given by a global vector field. This means that $\mathscr{F}\in{\rm Log}(1^{n-k+1},2;k,n)$, see Remark 2.2. ∎
Now, we apply this proposition to prove the general case.
Proof of Theorem 4.2.
Suppose that $\mathscr{F}$ is neither algebraically integrable nor logarithmic of type $(1^{n-k+1},2)$. By Proposition 4.3, we must have $h=h^{0}({\rm T}\mathscr{F}(-1))>0$. Then $\mathscr{F}$ is a linear pullback of a foliation $\mathscr{G}$ on $\mathbb{P}^{n-h}$} such that $H^{0}({\rm T}\mathscr{G}(-1))=0$. It follows that $\dim\mathscr{G}\leq 2$. Indeed, if we had $\dim\mathscr{G}\geq 3$ we would get a contradiction applying Proposition 4.3 again.
If $\dim\mathscr{G}=1$ then $\mathscr{F}\in{\rm LPB}(2,k,n)$. If $\dim\mathscr{G}=2$ then, by Theorem 4.1, $\mathscr{G}$ belongs to either ${\rm Log}(1^{n-h+1};2,n-h)$ or a Affine-Lie component ${\rm ALie}(r_{1},\dots,r_{m},2;n-h)$. Thus $\mathscr{F}$ belongs to either ${\rm Log}(1^{n+1};k,n)$ or ${\rm ALie}(r_{1},\dots,r_{m},k;n)$, see Remark 2.4. ∎
5. Holomorphic Poisson structures
In this section, we will prove Corollary C. A holomorphic Poisson structure on a complex manifold $X$ is a $\mathbb{C}$-linear Lie bracket
$\{\cdot,\cdot\}\colon\mathcal{O}_{X}\times\mathcal{O}_{X}\to\mathcal{O}_{X}$
which satisfies the Leibniz rule $\{f,gh\}=h\{f,g\}+f\{g,h\}$ and Jacobi identities
$\{f,\{g,h\}\}+\{g,\{h,f\}\}+\{h,\{f,g\}\}=0$
for all $f,g\in\mathcal{O}_{X}$. The bracket corresponds to a holomorphic bivector field $\sigma\in H^{0}(X,\wedge^{2}TX)$ given by $\sigma(df\wedge dg)=\{f,g\}$, for all $f,g\in\mathcal{O}_{X}$. We will denote a Poisson structure on $X$ as the pair $(X,\sigma)$, where $\sigma\in H^{0}(X,\wedge^{2}TX)$ is the corresponding Poisson bivector field. We say that the structure $(X,\sigma)$ has generic rank $2k$ if
$\underbrace{\sigma\wedge\cdots\wedge\sigma}_{k\ \mathrm{times}}\neq 0\quad\text{and}\quad\underbrace{\sigma\wedge\cdots\wedge\sigma}_{k+1\ \mathrm{times}}=0$.
The holomorphic bivector $\sigma$ induces a morphism
$\sigma^{\#}\colon\Omega_{X}^{1}\to TX$
which is called the anchor map and is defined by $\sigma^{\#}(\theta)=\sigma(\theta,\cdot)$, where $\theta$ is a germ of a holomorphic $1$-form, and $\delta\colon TX\to\wedge^{2}TX$ is given by $\delta(v)=L_{v}(\sigma)$; here $L_{v}$ denotes the Lie derivative. We refer to [34] for more details.
Definition 5.1.
The symplectic foliation associated with $\sigma$ is the foliation given by $\mathscr{F}_{\sigma}:=\mathrm{Ker}(\sigma^{\#})$, whose dimension is the rank of the anchor map $\sigma^{\#}:\Omega_{X}^{1}\to TX$. A Poisson variety $(X,\sigma)$ is called generically symplectic if the anchor map $\sigma^{\#}\colon\Omega_{X}^{1}\to TX$ is generically an isomorphism. Then, the degeneracy loci of $\sigma^{\#}$ is an effective anti-canonical divisor $D(\sigma)\in|-K_{X}|$.
Definition 5.2.
Let $(X,\sigma)$ be a Poisson projective variety. A Poisson connection on a sheaf of $\mathcal{O}_{X}$-modules $E$ is a $\mathbb{C}$-linear morphism of sheaves $\nabla:E\to TX\otimes E$ satisfying the Leibniz rule
$\nabla(fs)=\delta(f)\otimes s+f\nabla(s),$
where $f$ is a germ of a holomorphic function on $X$ and $s$ is a germ of a holomorphic section of $E$. We say that $E$ is a Poisson module if it admits a Poisson flat connection, i.e., if its curvature $\nabla^{2}\colon E\to\Omega_{X}^{2}\otimes E$ vanishes. Equivalently, a Poisson connection defines a $\mathbb{C}$-linear bracket $\{\ ,\ \}\colon\mathcal{O}_{X}\times E\to E$ by
$\{f,s\}:=\nabla(s)(df),$
where $f$ is a germ of a holomorphic function on $X$ and $s$ is a germ of a holomorphic section of $E$.
Let $E\cong\mathcal{O}_{X}^{\oplus 2}$ be a Poisson module with a trace free Poisson connection $\nabla$, then we can write
$\nabla=\delta+\begin{pmatrix}v_{1}&v_{2}\\ v_{0}&-v_{1}\end{pmatrix}$
and the flatness condition is equivalent to
$\begin{matrix}\delta(v_{0})=\hfill v_{0}\wedge v_{1}\\ \delta(v_{1})=2v_{0}\wedge v_{2}\\ \delta(v_{2})=\hfill v_{1}\wedge v_{2}\end{matrix}.$
We refer to [9] for more details.
Definition 5.3.
Let $D$ be a reduced divisor on $X$. The logarithmic tangent sheaf $TX(-\log(D))$ is the subsheaf of $TX$ consisting of germs of vector fields that preserve the ideal sheaf $\mathcal{O}_{X}(-D)$. That is, $TX(-\log(D))$ is the sheaf of vector fields tangent to $D$. A reduced divisor $D\subset X$ is free if $TX(-\log(D))$ is locally free. If $f:X\to Y$ is a rational map, we denote by $T_{X/Y}\subset TX$ the relative tangent sheaf and by $T_{X/Y}(-\log(D)):=T_{X/Y}\cap TX(-\log(D))$.
Definition 5.4.
An analytic subspace $Z\subset(X,\{\ ,\ \})$ is a Poisson subspace $Z$ is equipped with a Poisson structure $\{\ ,\ \}_{Z}$ such that the inclusion $i\colon Z\to X$ is a morphism of Poisson analytic spaces, i.e., it is compatible with the brackets.
Proposition 5.1.
[34, Proposition 4.4.1] Let $D\subset X$ a Poisson free divisor with respect to a holomorphic bi-vector $\sigma\in H^{0}(X,\wedge^{2}TX)$. Then $\sigma\in H^{0}(X,\wedge^{2}TX(-\log(D))$.
Proof of Corollary C.
Since the associated symplectic foliation $\mathscr{F}_{\sigma}$ has no divisorial zeros and $c_{1}({\rm T}\mathscr{F}_{\sigma})=0$, $\mathscr{F}_{\sigma}$ is a foliation of dimension $2$ and degree $2$ on ${\mathbb{P}}^{4}$. It follows from Theorem 4.1 that one of the following holds:
  1. (a)
    $\mathscr{F}_{\sigma}$ is algebraically integrable and $\sigma$ is not decomposable, since ${\rm T}\mathscr{F}_{\sigma}$ is stable.
  2. (b)
    $\mathscr{F}_{\sigma}$ is the linear pull-back of a degree-two foliation by curves on ${\mathbb{P}}^{3}$ with tangent sheaf split. So, $\sigma$ is decomposable.
  3. (c)
    ${\rm T}\mathscr{F}_{\sigma}\cong\mathfrak{g}\otimes\mathcal{O}_{{\mathbb{P}}^{4}}$, where $\mathfrak{g}\subset\mathfrak{sl}(5,\mathbb{C})$ and either $\mathfrak{g}$ is an abelian Lie algebra of dimension $2$ or $\mathfrak{g}\cong\mathfrak{aff}(\mathbb{C})$. So, $\sigma$ is decomposable.
  4. (d)
    $\mathscr{F}_{\sigma}$ is the pull-back by a rational map $\rho\colon{\mathbb{P}}^{4}\dashrightarrow\mathbb{P}(1^{(3)},2)$ of a non-algebraic foliation by curves induced by a global vector field on $\mathbb{P}(1^{(3)},2)$. Then ${\rm T}\mathscr{F}_{\sigma}$ is not split, which implies that $\sigma$ is not decomposable.
Case (a). Since $\sigma$ is of normal crossing type, by Remark 2.1 the symplectic foliation $\mathscr{F}_{\sigma}$ is tangent to a dominant rational map ${\mathbb{P}}^{4}\dashrightarrow{\mathbb{P}}^{2}$ given either by
$f_{(1,3)}:=(L_{1}^{3},L_{2}^{3},P)\quad\text{or}\quad f_{(1,2,2)}:=(L_{1}^{2},Q_{1},Q_{2}),$
where $L_{i}$ are linear, $Q_{i}$ are quadratic, and $P$ is cubic. Then the divisors
$D_{(1,3)}:=\{L_{1}L_{2}P=0\}\quad\text{and}\quad D_{(1,2,2)}:=\{L_{1}Q_{1}Q_{2}=0\}$
are Poisson divisors of $\sigma$. If we denote by $T_{{\mathbb{P}}^{4}/{\mathbb{P}}^{2}}$ the corresponding relative tangent sheaf either of $f_{(1,3)}\colon{\mathbb{P}}^{4}\dashrightarrow{\mathbb{P}}^{2}$ or of $f_{(1,2,2)}\colon{\mathbb{P}}^{4}\dashrightarrow{\mathbb{P}}^{2}$, then it follows from Proposition 5.1 that
  • $\sigma\in H^{0}({\mathbb{P}}^{4},\wedge^{2}T_{{\mathbb{P}}^{4}/{\mathbb{P}}^{2}}(-\log D_{(1,3)}))$, or
  • $\sigma\in H^{0}({\mathbb{P}}^{4},\wedge^{2}T_{{\mathbb{P}}^{4}/{\mathbb{P}}^{2}}(-\log D_{(1,2,2)}))$.
Case (b). If $\mathscr{F}_{\sigma}$ is the linear pull-back of a degree-two foliation by curves on ${\mathbb{P}}^{3}$, then there are $v_{0}$ a constant vector field and $v$ a homogeneous polynomial vector field of degree $2$ on ${\mathbb{P}}^{3}$ such that $\sigma=v_{0}\wedge v$. It is clear that $v_{0},v]=0$. Let $H$ be the Poisson hyperplane determined by $v_{0}\in\mathbb{P}H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}(-1))\cong\check{{\mathbb{P}}}^{4}$. That is, $H$ is the hyperplane parameterizing lines passing through a point which are leaves of the foliation induced by $v_{0}$. It is clear that $H$ is a Poisson divisor.
Case (c). If ${\rm T}\mathscr{F}_{\sigma}\cong\mathfrak{g}\otimes\mathcal{O}_{{\mathbb{P}}^{4}}$, where $\mathfrak{g}\subset\mathfrak{sl}(5,\mathbb{C})$ is an abelian Lie algebra, then since $\mathscr{F}_{\sigma}$ is general, the Lie algebra $\mathfrak{g}$ is generated by diagonal linear vector fields which are given, in a suitable system of homogeneous coordinates $z_{0}:\dots:z_{4}]$ on ${\mathbb{P}}^{4}$, by
$v_{0}=\sum_{i=0}^{4}\lambda_{0i}z_{i}\dfrac{\partial}{\partial z_{i}}\quad\text{and}\quad v_{1}=\sum_{j=0}^{4}\lambda_{1j}z_{j}\dfrac{\partial}{\partial z_{j}}.$
Then
$\sigma=v_{0}\wedge v_{1}=\sum_{i<j}(\lambda_{0i}-\lambda_{1j})\,z_{i}z_{j}\,\dfrac{\partial}{\partial z_{i}}\wedge\dfrac{\partial}{\partial z_{j}}.$
Since $v_{r}(z_{i})=\lambda_{ri}z_{i}$ for all $r=0,1$ and $i=0,\dots,4$, we conclude that $v_{0},v_{1}\in H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}(-\log D))$ and $\sigma\in\wedge^{2}\mathfrak{h}$, where
$\mathfrak{h}=H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}(-\log D))$
is the abelian Lie algebra of vector fields tangent to the arrangement of coordinate hyperplanes
$D=\{z_{0}z_{1}z_{2}z_{3}z_{4}=0\}.$
Suppose now that ${\rm T}\mathscr{F}_{\sigma}\cong\mathfrak{g}\otimes\mathcal{O}_{{\mathbb{P}}^{4}}$, where $\mathfrak{g}\cong\mathfrak{aff}(\mathbb{C})$. In this case, we have that the Poisson structure on ${\mathbb{P}}^{4}$ is induced by the so-called infinitesimal action
$\sigma\colon\mathfrak{g}\otimes\mathcal{O}_{{\mathbb{P}}^{4}}\longrightarrow T{\mathbb{P}}^{4},$
which is generated by linear vector fields $v_{0}$ and $v_{1}$ such that $v_{0},v_{1}]=v_{0}$. In addition, in both cases where $\sigma=v_{0}\wedge v_{1}$, with $v_{0},v_{1}\in H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4})$, we have that the tangent sheaf of the symplectic foliation is ${\rm T}\mathscr{F}_{\sigma}\cong\mathcal{O}_{{\mathbb{P}}^{4}}^{\oplus 2}$. We have a trace-free flat Poisson connection
$\nabla\colon{\rm T}\mathscr{F}_{\sigma}\longrightarrow T{\mathbb{P}}^{4}\otimes{\rm T}\mathscr{F}_{\sigma}$
given by
$\nabla=\delta+\begin{pmatrix}v_{1}&0\\ v_{0}&-v_{1}\end{pmatrix}$.
If ${\rm T}\mathscr{F}_{\sigma}\cong\mathfrak{g}\otimes\mathcal{O}_{{\mathbb{P}}^{4}}$ with $\mathfrak{g}$ abelian, then for $i=0,1$ we have
$\delta(v_{i})=L_{v_{i}}(v_{0}\wedge v_{1})=0,$
since $v_{1},v_{0}]=0$. If ${\rm T}\mathscr{F}_{\sigma}\cong\mathfrak{g}\otimes\mathcal{O}_{{\mathbb{P}}^{4}}$ with $\mathfrak{g}\cong\mathfrak{aff}(\mathbb{C})$, then
$\delta(v_{0})$ $=L_{v_{0}}(v_{0}\wedge v_{1})=v_{0}\wedge v_{1},$
$\delta(v_{1})$ $=L_{v_{1}}(v_{0}\wedge v_{1})=0,$
since $v_{1},v_{0}]=-v_{1}$. Now, let $p\in{\mathbb{P}}^{4}\setminus Sing(\mathscr{F}_{\sigma})$ and let $L_{p}$ be the symplectic leaf of $\mathscr{F}_{\sigma}$ passing through $p$. Since
$(\sigma^{\#})^{-1}\circ\delta|_{L_{p}}\;=\;d\circ(\sigma^{\#})^{-1}|_{L_{p}},$
where $d$ denotes the de Rham differential, we have that
$\begin{matrix}d\circ(\sigma^{\#})^{-1}(v_{0})=(\sigma^{\#})^{-1}\circ\delta|_{L_{p}}(v_{0})=(\sigma^{\#})^{-1}(v_{0})\wedge(\sigma^{\#})^{-1}(v_{1})\\[4.0pt] d\circ(\sigma^{\#})^{-1}(v_{1})=(\sigma^{\#})^{-1}\circ\delta|_{L_{p}}(v_{1})=2(\sigma^{\#})^{-1}(v_{0})\wedge(\sigma^{\#})^{-1}(v_{2}).\end{matrix}$
Now, defining $(\sigma^{\#})^{-1}|_{L_{p}}(v_{i})=\omega_{i}$, we obtain the transversely affine structure on $L_{p}$ given by the pair $(\omega_{0},\omega_{1})$, where $\omega_{0}$ is the $1$-form inducing the transversely affine holomorphic foliation on $L_{p}$. We have:
  • if $\mathfrak{g}$ is abelian, then $(\omega_{0},\omega_{1})$ is Euclidean;
  • if $\mathfrak{g}\cong\mathfrak{aff}(\mathbb{C})$, then $(\omega_{0},\omega_{1})$ is affine.
If ${\rm T}\mathscr{F}_{\sigma}\cong\mathfrak{g}\otimes\mathcal{O}_{{\mathbb{P}}^{4}}$, where $\mathfrak{g}\cong\mathfrak{aff}(\mathbb{C})$, then the normal form of $\sigma$ is given by the vector fields in (2.5) specialized to $n=4$.
Case (d). The symplectic foliation $\mathscr{F}_{\sigma}$ is the pull-back by a rational map
$\rho\colon{\mathbb{P}}^{4}\dashrightarrow\mathbb{P}(1^{(3)},2)$
of a foliation by curves on $\mathbb{P}(1^{(3)},2)$, and
$\rho:=(L_{1},L_{2},L_{3},Q),$
where $L_{i}$ are linear and $Q$ is quadratic. Then
$\sigma\in H^{0}({\mathbb{P}}^{4},\wedge^{2}T{\mathbb{P}}^{4}(-\log D)),\quad D:=\{L_{1}L_{2}L_{3}Q=0\}.$
Consider
$v\in H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}(-\log D))\subset H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4}),$
the global vector field whose leaves are fibers of $\rho\colon{\mathbb{P}}^{4}\dashrightarrow\mathbb{P}(1^{(3)},2)$. It is clear that $\sigma\wedge v=0$ and $dF(v)=0$, where $F:=L_{1}L_{2}L_{3}\cdot Q$. That is, $v$ is Hamiltonian with respect to $\sigma$.
Finally, in the case $\sigma=v_{0}\wedge v_{1}\in\wedge^{2}\mathfrak{h}$, where $\mathfrak{h}=H^{0}({\mathbb{P}}^{4},T{\mathbb{P}}^{4})$ and $v_{0}$ and $v_{1}$ are linear vector fields such that $v_{0},v_{1}]=v_{0}$, the normal form follows from subsection 2.2.1. ∎
Remark 5.1.
Let $({\mathbb{P}}^{4},\sigma)$ be a holomorphic Poisson structure such that $\sigma\wedge\sigma=0$. Then the associated symplectic foliation $\mathscr{F}_{\sigma}$ has degree at most $2$. If $\sigma$ has zeros of codimension one, then by the classification of foliations of degree 0 and $1$ (see Section 2.1), we have that either
  • $\sigma=Q\cdot(v_{0}\wedge v_{1})$, where $v_{i}$ are constant vector fields and $Q$ is of degree $2$, or
  • $\sigma=H_{0}\cdot\tilde{\sigma}$, with $\{H_{0}=0\}$ being a Poisson hyperplane and
    • $\tilde{\sigma}=v_{0}\wedge v$, where $v_{0}$ is a constant vector field, $v$ is a linear vector field such that $v_{0},v]=0$;
    • $\tilde{\sigma}\in H^{0}({\mathbb{P}}^{4},\wedge^{2}T_{{\mathbb{P}}^{4}/{\mathbb{P}}^{2}}(-\log D))$, where $D=H_{1}\cup H_{2}\cup H_{3}\cup\mathcal{Q}$ is a Poisson divisor such that $H_{i}$ are linear and $\mathcal{Q}$ is of degree $2$.